### v1

# CAMPBELL & REMIRO-AZÓCAR

Abstract
Externally controlled single-arm trials are critical to assess treatment efficacy across therapeutic indications for which randomized controlled trials are not feasible. A closely-related research design, the unanchored indirect treatment comparison, is often required for disconnected treatment networks in health technology assessment. We present a unified causal inference framework for both research designs. We develop a novel estimator that augments a popular weighting approach based on entropy balancing – matching-adjusted indirect comparison (MAIC) – by fitting a model for the conditional outcome expectation. The predictions of the outcome model are combined with the entropy balancing MAIC weights. While the standard MAIC estimator is singly robust where the outcome model is non-linear, our augmented MAIC approach is doubly robust, providing increased robustness against model misspecification. This is demonstrated in a simulation study with binary outcomes and a logistic outcome model, where the augmented estimator demonstrates its doubly robust property, while exhibiting higher precision than all non-augmented weighting estimators and near-identical precision to G-computation. We describe the extension of our estimator to the setting with unavailable individual participant data for the external control, illustrating it through an applied example. Our findings reinforce the understanding that entropy balancing-based approaches have desirable properties compared to standard “modeling” approaches to weighting, but should be augmented to improve protection against bias and guarantee double robustness.
keywords:
Single-arm trial, indirect treatment comparison, external control, covariate adjustment, evidence synthesis, data fusion
1 Background
In pharmaceutical research, randomized controlled trials (RCTs) are the gold standard for evaluating the efficacy and safety of new treatments, due to their high internal validity. The random allocation of subjects to either an experimental therapy or a concurrent control minimizes the risk of confounding, balancing in expectation the distribution of prognostic factors over the treatment arms. Nevertheless, conducting an RCT is not possible in certain settings. For instance:
  • Where recruitment to properly-powered RCTs is unfeasible or impractical due to small populations and a lack of trial-eligible patients, e.g., rare diseases with orphan designation or biomarker-specific precision oncology indications;1
  • For life-threatening and severely debilitating conditions with high unmet medical need, due to non-existent or inadequate standard of care, e.g., “last-line of therapy” indications in late-stage hematological and solid tumor oncology;2
  • Where the enrollment of patients to a placebo or “no treatment” control arm is unethical, e.g., withholding a therapy with proven efficacy in adults from a control group of children in pediatric trials.3
Regulatory agencies such as the Food and Drug Administration (FDA) and the European Medicines Agency (EMA) emphasize that properly conducted RCTs provide the highest evidentiary standard for the approval of new drugs.4, 5 Nevertheless, regulators recognize that alternative research designs may be required in special circumstances. One of such designs is the externally controlled single-arm trial (SAT), where the control group is fully derived from external data such as prior clinical trials or secondary real-world data (RWD) sources, e.g., electronic health records, disease registries or medical claims.2, 6, 7, 8
Marketing authorization applications featuring externally controlled SATs continue to rise, especially under the accelerated approval pathway from the FDA and the conditional approval pathway from the EMA.8, 9 In 2023 the FDA issued a draft guidance document for externally controlled trials,10 and in 2024 the EMA finalized a reflection paper on single-arm trials.11 Regulators have granted approvals to new therapeutics based on externally controlled SATs, especially in rare diseases, oncology and hematology, with approval being more likely for conditions with a highly predictable natural history, precisely measurable endpoints, and where a large effect size is anticipated.1, 2, 8, 9, 12, 13
For health technology assessment (HTA), head-to-head RCT evidence also remains the gold standard for decision-making.14, 15, 16 Nevertheless, as regulatory authorities increasingly approve pharmaceuticals based on externally controlled SATs, the reliance of payers on such research designs has grown.9, 17 HTA bodies are developing recommendations for externally controlled SATs in reimbursement submissions,18, 19 with acceptability being greater in oncology and also influenced by factors such as high unmet need and the rarity of the disease.9, 17, 20
Additionally, HTA requires comparing the clinical effectiveness and cost-effectiveness of new health technologies against all existing alternatives on the market.21 The scope of assessments often depends on the policy question and is not always driven by the available data.22 A single RCT cannot typically have all the treatment arms that are desired for HTA, particularly given the multiplicity of national and regional stakeholders, and the variations in clinical practice across jurisdictions, which may themselves arise due to historical variations in reimbursement practice.23 In the absence of direct RCT comparisons versus all candidate comparators, indirect treatment comparisons (ITCs) across studies are required.24
HTA decision-makers have expressed a clear preference for anchored ITCs of randomized trials.25, 26, 27 These respect study randomization by using a common control arm to contrast relative treatment effects across trials.28 Nevertheless, even for indications where RCTs are regularly conducted, at times it is not possible to find a compatible control arm with which to “anchor” the analysis. This is especially the case in rapidly evolving therapeutic areas with multiple novel treatments entering the market, a changing comparator landscape and no single accepted standard of care internationally.29, 30 In these scenarios, unanchored ITCs based on disconnected treatment networks may be required29, 30 and a recent literature review determined that unanchored ITCs are in fact far more common than anchored ITCs. 31
Unanchored ITCs contrast mean treatment-specific absolute outcomes across studies, thereby relying on more restrictive assumptions than anchored ITCs.25, 28 In essence, unanchored ITCs are externally controlled SATs with two specific characteristics. Firstly, the external control is often a historical trial that has been sponsored by a competitor. Secondly, due to privacy and confidentiality reasons, there is limited access to subject-level data. While individual participant data (IPD) are available for the SAT, only published aggregate-level data (AD) are available for the external control.25, 28
The absence of randomization compromises the validity of externally controlled SATs. Various statistical methods have been proposed to adjust for imbalances in baseline covariates between the SAT and the external control.32 These can potentially mitigate confounding bias and account for the additional variability induced by covariate differences. The most widely-used techniques are propensity score-based weighting approaches, which explicitly model the conditional probability of SAT participation – a propensity score – as a function of baseline covariates. Typically, a logistic regression is applied, such that the logit of the propensity score is assumed to vary linearly with the covariates, and this is estimated via maximum-likelihood.33, 34, 35, 36
Specifically within the context of unanchored ITCs, an alternative weighting method based on entropy balancing37 called matching-adjusted indirect comparison (MAIC) is more popular.38, 39, 40 This views covariate balance as a convex optimization problem, estimating weights that directly enforce balance in covariate moments between the SAT and the external control, without explicitly modeling the conditional probability of SAT participation. MAIC is attractive for ITCs due to its applicability in the “IPD-AD” situation, where there is limited access to IPD for the external control. Moreover, entropy balancing techniques such as MAIC are thought to be generally more stable, more precise, and more robust to model misspecification relative to the standard “modeling” approaches to weighting, even in “IPD-IPD” scenarios.40, 41, 42, 43
So-called “G-computation” or “model-based standardization” methods have also been developed for the IPD-IPD33, 34, 44, 45 and IPD-AD settings.46 For the latter, they are also referred to as “simulated treatment comparison” (STC).47 These methods are based on estimating a model for the conditional outcome expectation given baseline covariates (the “outcome model”), and averaging model-based outcome predictions over the target covariate distribution. G-computation approaches gain statistical precision and efficiency with respect to weighting, particularly where poor covariate overlap leads to extreme weights and large reductions in effective sample size.46 That being said, the improved performance relies on the extrapolation of a correctly specified outcome model and G-computation methods can be prone to bias and undercoverage, even under relatively minor model misspecification.48
Weighting and G-computation techniques make different modeling assumptions, but both are generally “singly robust”. The former, in most cases, depends on correct specification of the propensity score model for consistent – and therefore asymptotically unbiased – estimation. The latter depends on correct specification of the model for the conditional outcome expectation. Decision-makers have expressed a preference for “doubly robust” estimation approaches that can consistently estimate the treatment effect as long as either the propensity score model or the outcome model is correct, but not necessarily both.25, 27, 28, 49 These methods should reduce the risk of bias by offering two opportunities for correct model specification. Despite this, doubly robust methods have rarely been applied in the analysis of externally controlled SATs.50 To the best of our knowledge, they have never been applied to unanchored ITCs, despite their development being recommended by HTA agencies.25, 28
One reason why doubly robust methods do not feature prominently for unanchored ITCs might be a misunderstanding that MAIC, the most popular approach for ITCs,31 is always doubly robust. To be clear, MAIC does enable consistent estimation when an implicit propensity score model is misspecified, but only if the true outcome model is linear with respect to the covariate functions that are balanced. As such, entropy balancing is said to be “linearly doubly robust”,40, 42, 43 with Zhao and Percival (2017) stating that it is “doubly robust with respect to linear outcome regression and logistic propensity score regression”.43 In practice, it is rarely plausible that outcomes vary linearly with the covariates (e.g., due to non-linear link functions or outcomes that depend on non-linear transformations of the covariates). Doubly robust methods for ITCs that are not necessarily restricted to linear outcome models are yet to be developed, with Josey et al (2021) recently identifying this as a research priority.40
The objective of this paper is to clarify existing approaches for doubly robust estimation in the analysis of externally controlled SATs and propose a doubly robust augmented MAIC estimator, specifically tailored to unanchored ITCs. In Section 2, we introduce and motivate the target estimand. In Section 3, we outline the available singly and doubly robust estimators and our proposed doubly robust estimator for unanchored ITCs, explaining the rationale behind the latter being doubly robust. In Section 4, we empirically evaluate the statistical performance of the approaches in a simulation study, comparing doubly robust estimators against the singly robust estimators and other augmented estimators that have also been described as doubly robust, such as G-computation with an outcome model fitted using weighted maximum-likelihood estimation.51, 52 The simulations demonstrate proof-of-principle with binary outcomes, a logistic outcome model and the marginal log-odds ratio as the target summary effect measure, where entropy balancing approaches do not have the doubly robust property. In Section 5, we illustrate the application of the methods in an example analysis. Lastly, we discuss our findings and conclude the paper in Section 6.
2 Estimands
Prior to introducing the statistical methodology, we define the different estimands that can be targeted by externally controlled SATs. An estimand is a precise definition of the treatment effect, which should align with the clinical question of interest, the research design and the analytical approach. Within the regulatory environment, the use of estimands has been stimulated by the publication of the International Council of Harmonisation E9 (R1) Addendum, adopted by the FDA and the EMA. While E9 (R1) was originally proposed for registrational RCTs, its principles are applicable to externally controlled SATs.53 According to E9 (R1), an estimand encompasses five attributes: target population, treatment(s), endpoint (outcome), summary effect measure and the strategy for handling intercurrent events.
We will focus specifically on the “target population” and “summary effect measure” components when defining the estimands, which are:
  • The average treatment effect (ATE) among the combined SAT and external control;
  • The average treatment effect in the treated (ATT); that is, among those participating in the SAT;
  • The average treatment effect in the control (ATC); that is, among those in the external control group.
The difference between these summary effect measures is driven by them targeting different (sub) populations or applying to different “analysis sets”. Having assumed that the SAT and external control analysis sets are random samples of their underlying target populations, we will generally make no further distinction between sample-level and population-level estimands.
We adopt the potential outcomes notation to define the estimands more formally. Let $Y^{t}$ represent the potential outcome that would have been observed for a subject under intervention $T=t$, with $t\in\{0,1\}$, where $T=1$ denotes an active intervention under investigation in the SAT ($S=1$), and $T=0$ denotes a pertinent control for the subjects in the external data source ($S=0$). Two potential outcomes, $(Y^{1},Y^{0})$, are defined for every subject, regardless of whether they are members of the SAT under $T=1$ or of the external control under $T=0$. Assuming no loss to follow up, one of the potential outcomes is observed for a given subject and the other is the “counterfactual” which would be realized under a different treatment than that actually assigned.
The ATE is defined as:
$\textrm{ATE}=g\left(\textrm{E}\left(Y^{1}\right)\right)-g\left(\textrm{E}\left(Y^{0}\right)\right),$
on the additive scale imposed by link function $g(\cdot)$, which transforms the potential outcome means into the plus/minus infinity range. Each expectation, $\textrm{E}(\cdot)$, is taken over the distribution of potential outcomes in the combined SAT and external control population. For a binary outcome, suitable link functions could be the identity, log or logit, to produce a risk difference, log relative risk or log-odds ratio as the summary effect measure. The ATT is defined as:
$\textrm{ATT}=g\left(\textrm{E}(Y^{1}\mid S=1)\right)-g\left(\textrm{E}\left(Y^{0}\mid S=1\right)\right),$
such that expectations are taken over the distribution of potential outcomes in the SAT (sub) population. The ATC is defined as:
$\textrm{ATC}=g\left(\textrm{E}(Y^{1}\mid S=0)\right)-g\left(\textrm{E}(Y^{0}\mid S=0)\right),$
such that expectations are taken over the distribution of potential outcomes in the external control (sub) population.
Within RCTs, the ATE, ATT and ATC are identical in expectation. However, they generally differ in externally controlled SATs, and will almost invariably do so where there is (conditional) treatment effect heterogeneity by the covariates, i.e., effect measure modification. We view the ATE target population, defined by pooling the SAT and the external control, as somewhat ambiguous in this context. As such, the target estimand in an externally controlled SAT is often either the ATT or the ATC.
The ATT is typically the primary estimand for drug approval purposes in the regulatory environment. It is consistent with emulating a randomized comparison in the pivotal trial population, with the external control mimicking the internal comparator arm of a registrational clinical trial. The ATT is also compatible with the mean absolute outcome that is targeted by the SAT, $\textrm{E}(Y^{1}\mid S=1)$, preserving the original SAT results. Nevertheless, SAT populations are often highly selected and may lack representativeness with respect to “real-world” patient populations. As such, the ATT may be less appealing where generalizability to routine clinical practice is a priority, as is the case in HTA decision-making.
The ATC is potentially more desirable for external validity. External controls based on natural history studies and RWD have broad inclusion criteria and populations that are relatively heterogeneous in clinical and demographic characteristics. However, external controls are not always representative of routine clinical practice. Historical controls from past clinical trials will not reflect the current standard of care due to the passage of time. Moreover, RWD-derived external controls are often based on a single country and not necessarily transferable to the relevant jurisdiction for decision-making.
Sample size considerations and challenges in statistical estimation may also play a part in the estimand choice. Both SATs and external controls often have low sample sizes, particularly in rare diseases. Covariate adjustment may reduce effective sample sizes further. The ATT or ATC may not be reliably estimated where sample size is low in the external control or the SAT, respectively. Consider weighting, where the estimand impacts the definition of the weights. Targeting the ATT implies preserving the original SAT, re-weighting and reducing the effective sample size of the external control. Conversely, targeting the ATC implies preserving the original external control, re-weighting and reducing the effective sample size of the SAT.
Finally, where IPD are available for the SAT but only published AD are available for the external control, as is typically the case in unanchored ITCs, the target estimand is often the ATC by necessity.25, 28, 32 Throughout the rest of the manuscript, we shall assume that there is unlimited access to subject-level data but that the ATC is the primary target of estimation. Our methodological approaches and findings are also applicable to the ATT, or where subject-level data are unavailable for the external control, with some caveats that will be discussed in Section 3.
3 Methodology
3.1 Data and assumptions
As per Section 2, let $T=t$ denote a time-invariant binary treatment, with $t\in\{0,1\}$, such that $T=1$ represents the active intervention and $T=0$ the control. Let $S=s$ denote the data source, with $s\in\{0,1\}$, such that $S=1$ represents the SAT and $S=0$ the external data source. In addition, let $\mathbf{X}$ denote vector-valued pre-treatment baseline covariates, e.g., clinical or demographic characteristics, measured across the SAT and the external data source. Let $Y$ denote the clinical outcome of interest. We assume that only distributional differences in $\mathbf{X}$ are preventing exchangeability between the SAT and external subjects, and that covariates and outcomes are defined and measured similarly across data sources.
The observed IPD consist of $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$, $i=1,\dots,n=n_{1}+n_{0}$, realizations of $(S,\mathbf{X},T,Y)$ denoting the data source, baseline covariates, treatment assignment and observed outcome for subject $i$. Here, the SAT and external data source have been stacked, with $n_{1}$ and $n_{0}$ as the sample sizes of the SAT and the external data source, respectively. It is assumed that all individuals in the SAT are under $T=1$ and all individuals in the external data source are under $T=0$, such that the control group is fully external. To be clear, we have $S_{i}=T_{i}$ for all $i=1,\dots,n$; and also have that $S_{i}=1$ and $T_{i}=1$ for all $i=1,\dots,n_{1}$, and $S_{i}=0$ and $T_{i}=0$ for all $i=n_{1}+1,\dots,n$. We shall assume that there is no missingness or measurement error.
The observed outcome for subject $i$ is $Y_{i}=Y_{i}^{1}T_{i}+Y_{i}^{0}(1-T_{i})$, where $Y_{i}^{t}$ is the potential outcome had subject $i$ been assigned treatment $t\in\{0,1\}$, with $Y_{i}=Y_{i}^{1}$ if $i=1,\dots,n_{1}$ and $Y_{i}=Y_{i}^{0}$ if $i=n_{1}+1,\dots,n$. Namely, the observed outcome for an individual in the SAT equals their potential outcome under the active intervention, and the observed outcome for an individual in the external data source equals their potential outcome under the control. Implicit in the notation is the stable unit treatment value assumption (SUTVA): that there is no interference between subjects and there is treatment version irrelevance, i.e., one well-defined version of the active intervention and the control across all subjects and data sources.54 Also implicit is that there is no direct effect of trial participation.54 Namely, that trial participation – in the SAT or a historical trial, for that matter – does not affect the outcome except through treatment assignment itself, i.e., there are no Hawthorne effects.55, 56
To estimate the ATC, we must construct estimators for $\mu^{1}_{0}=\textrm{E}(Y^{1}\mid S=0)$ and $\mu^{0}_{0}=\textrm{E}(Y^{0}\mid S=0)$. Outcomes for the subjects from the external data source have been generated under the control and we assume that there is no informative missingness or measurement error. Hence, unbiased estimation of $\mu^{0}_{0}$ should be trivial using the sample mean, such that $\hat{\mu}^{0}_{0}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}$. Conversely, while the active intervention has been investigated in the SAT, its outcomes in the external control (sub) population are unobserved. Our challenge is therefore to produce a reliable estimate $\hat{\mu}^{1}_{0}$ of the mean absolute outcome $\mu^{1}_{0}$ under the active intervention in the external control (sub) population, based on the observed data.
[…]
To enforce positivity, analysts may subset the SAT based on the selection criteria of the external control.25, 28 However, this further reduces the sample size of the SAT. Positivity is typically assessed by comparing the empirical distributions of the covariates in the SAT and the external control.60 While outcome modeling-based approaches such as G-computation can overcome failures of positivity, they do so by potentially problematic and difficult-to-diagnose model-based extrapolation. Even minor model misspecification over the observed covariate space in the SAT may lead to poor extrapolation in unobserved regions.48
Analogously, targeting the ATT would require constructing estimators for $\mu^{1}_{1}=\textrm{E}(Y^{1}\mid S=1)$ and $\mu^{0}_{1}=\textrm{E}(Y^{0}\mid S=1)$. Here, the challenge is the estimation of $\mu^{0}_{1}$ because outcomes under the control have not been generated in the SAT. The conditional ignorability assumption would formally be $Y_{i}^{0}\perp\!\!\!\!\perp S_{i}\mid\mathbf{x}_{i}$ for all $i=1,\dots,n$, and would rest on the SAT and the external control measuring all variables that are prognostic of outcome under the control. The positivity assumption would be $10<\textrm{Pr}(S=0\mid\mathbf{X}=\mathbf{x})<$ 1 for all $\mathbf{x}$ with positive density in the SAT, $f(\mathbf{x}\mid S=1)>0$, such that the support of the baseline covariates in the SAT is contained within that of the external control and it is possible to have external control subjects in all regions of the SAT covariate distribution. In this setting, analysts may apply the SAT selection criteria to the external control to guarantee that there is sufficient overlap.
3.2 Inverse odds weighting
We first present a covariate adjustment method that models the data source assignment mechanism, conditional on baseline covariates, to estimate weights.33, 34, 40, 61, 62 Where the target estimand is the ATC, SAT subjects are weighted by their inverse conditional odds of SAT participation – their conditional odds of external control participation – to transport the SAT outcomes to the external control (sub) population. Such “inverse odds” weights (IOW) are defined as:
$w_{i}=\frac{(1-e_{i})S_{i}}{e_{i}}+(1-S_{i}),$ (1)
for subject $i=1,\dots,n$, where the propensity score $e_{i}=e(\mathbf{X}_{i})=\textrm{Pr}(S_{i}=1\mid\mathbf{X}_{i})$ denotes the conditional probability of SAT participation given covariates $\mathbf{X}_{i}$ for subject $i$. In Equation 1, note that the SAT subjects $(S_{i}=1)$ are weighted as $w_{i}=(1-e_{i})/e_{i}$, whereas the external control subjects $(S_{i}=0)$ are unweighted, i.e., assigned a weight of $w_{i}=1$.
In practice, the true propensity scores are unknown. Almost invariably, there are multiple baseline covariates and at least one of these is continuous, such that a data source assignment model is required to estimate the propensity scores. The model is often a logistic regression:
$\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\alpha},$ (2)
where $\operatorname{logit}(e_{i})=\ln\left((e_{i})/(1-e_{i})\right)$, $\alpha_{0}=\ln\left(\textrm{Pr}(S_{i}=1|\mathbf{c}(\mathbf{X}_{i})=0)/(\textrm{Pr}(S_{i}=0|\mathbf{c}(\mathbf{X}_{i})=0)\right)$ is an intercept term, $\mathbf{\alpha}$ is a vector of regression parameters, and $\mathbf{c}(\mathbf{X}_{i})=[c_{1}(\mathbf{X}_{i}),c_{2}(\mathbf{X}_{i}),\dots,c_{k}(\mathbf{X}_{i})]^{\top}$ is a vector of covariate “balance functions” for subject $i=1,\dots n$. This is the set of functions containing the distributional features to be balanced between the SAT and the external control,40, 63 potentially including sensible transformations of the covariates, e.g., polynomials and interaction terms.
The logistic regression is typically fitted to the concatenated IPD using maximum-likelihood estimation, with the regression coefficient point estimates denoted by $\hat{\alpha}_{0}$ and $\hat{\mathbf{\alpha}}$, and model-based propensity scores for subject $i=1,\dots,n_{1}$ predicted by $\hat{e}_{i}=\operatorname{logit}^{-1}\left(\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\alpha}}\right)=\operatorname{expit}\left(\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\alpha}}\right)$, where $\operatorname{expit}(\cdot)=\exp(\cdot)/\left(1+\exp(\cdot)\right)$. Weight estimates $\hat{w}_{i}$ for $i=1,\dots,n_{1}$ are derived by plugging the corresponding propensity score predictions into Equation 1. With correct specification of the model in Equation 2, such that the log-odds of SAT participation are linear across the balance functions of the covariates, $\hat{e}_{i}$ and $\hat{w}_{i}$ consistently estimate the true conditional probability and inverse odds of SAT participation, respectively.
The ATC is estimated by contrasting the weighted average of observed outcomes under the active intervention with the unweighted average of observed outcomes for the external control. As per Section 2, mean absolute outcomes are converted to the additive scale imposed by link function $g(\cdot)$ prior to taking the difference between treatments on such scale:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=1}^{n_{1}}\hat{w}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (3)
The mean absolute outcome estimate for the active intervention can be bounded within its feasible range, e.g., between 0 and 1 for probabilities, by normalizing or “stabilizing” the weights so that they sum to one.57, 61 This results in the alternative ATC estimator:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{\sum_{i=1}^{n_{1}}\hat{w}_{i}Y_{i}}{\sum_{i=1}^{n_{1}}\hat{w}_{i}}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (4)
which should provide improved finite sample properties and more stable and precise estimation.64, 65
In expectation, if the model in Equation 2 is correctly specified, the estimated weights, ($\hat{w}_{i}$, for $i$ in $n_{1}$) will balance the covariate distribution of the SAT with respect to that of the external control, enabling consistent estimation of mean absolute outcome $\mu^{1}_{0}$ and the ATC. To see that the IOW estimators are consistent, consider a simple scenario with a binary outcome, $Y$, and a single discrete covariate, $X$, such that:
$\mu_{0}^{1}=\textrm{E}(Y|T=1,S=0)$ $=1\times\textrm{Pr}(Y=1|T=1,S=0)+{0\times\textrm{Pr}(Y=0|T=1,S=0)}$
$=\textrm{Pr}(Y=1|T=1,S=0).$
Basic probability rules imply that the marginal risk is the weighted average of the stratum-specific risks:
$=\sum_{x}\textrm{Pr}(Y=1|T=1,S=0,X=x)\textrm{Pr}(X=x|S=0,T=1).$
Then, due to the assumption of conditional data source ignorability, we have:
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=0)$
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=0)\frac{\textrm{Pr}(X=x|S=1)}{\textrm{Pr}(X=x|S=1)}$
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=1)\frac{\textrm{Pr}(X=x|S=0)}{\textrm{Pr}(X=x|S=1)},$
and from Bayes’ rule, we have:
$=\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=1)\frac{\textrm{Pr}(S=0|X=x)}{\textrm{Pr}(S=1|X=x)}\frac{\textrm{Pr}(S=1)}{\textrm{Pr}(S=0)}$
$=\underbrace{\frac{\textrm{Pr}(S=1)}{\textrm{Pr}(S=0)}}_{A}\underbrace{\sum_{x}\textrm{Pr}(Y=1|T=1,X=x)\textrm{Pr}(X=x|S=1)\frac{\textrm{Pr}(S=0|X=x)}{\textrm{Pr}(S=1|X=x)}}_{B}.$ (5)
Since $\sum_{i=1}^{n}{\frac{S_{i}}{n}}=n_{1}/n\rightarrow\textrm{Pr}(S=1)$ and $\sum_{i=1}^{n}{(1-S_{i})}/n=n_{0}/n\rightarrow\textrm{Pr}(S=0)$, we can consistently estimate $A\approx{n_{1}}/{n_{0}}$. We can also consistently estimate $B$ from the sample since the covariate distribution for the SAT, $(X|S=1)$, is observed (and since, within the sample, $T=1\iff S=1$):
$B$ $\approx\sum_{i=1}^{n_{1}}\textrm{Pr}(Y=1|S=1,X=x_{i})\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}\approx\sum_{i=1}^{n_{1}}\frac{Y_{i}}{n_{1}}\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}.$
Finally, if the propensity score model is correctly specified, the inverse odds weights consistently estimate the true inverse odds, such that, for $i$ in $n_{1}$:
$\widehat{w}_{i}\rightarrow$ $\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}.$ (6)
Therefore, the IOW weighted estimator is consistent:
$\hat{\mu}_{0}^{1}$ $=\underbrace{\frac{n_{1}}{n_{0}}}_{A}\underbrace{\sum_{i=1}^{n_{1}}{\frac{Y_{i}}{n_{1}}\hat{w}_{i}}}_{B}$
$=\frac{1}{n_{0}}\sum_{i=1}^{n_{1}}{Y_{i}\hat{w}_{i}}\rightarrow\mu_{0}^{1}.$ (7)
Note that $\textrm{E}(S_{i}w_{i})=n_{0}/n_{1}$ implies that $\textrm{E}\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}{w_{i}}\right)=n_{0}/n_{1}$, which implies that $\textrm{E}\left(\sum_{i=1}^{n_{1}}{w_{i}}\right)=n_{0}$. As such, the stabilized IOW estimator in Equation 4 is also consistent for the ATC.
3.3 Entropy balancing (matching-adjusted indirect comparison)
In Section 3.2, we described a “modeling approach” to weighting.66, 67 Propensity scores are explicitly modeled as a function of baseline covariates by a logistic regression, and are estimated by maximizing the fit of such regression, with inverse odds weights derived from the estimated propensity scores. This approach has certain limitations:66, 67
  • The estimated weights do not produce adequate covariate balance if the propensity score model is misspecified, and even a correctly specified model does not guarantee balance in finite samples;
  • Propensity score predictions that are close to zero produce extreme and highly variable weights, which lead to unstable and imprecise ATC estimation, particularly where overlap is poor or the sample size of the SAT is small; and
  • There is limited applicability when covariate IPD for the external control are unavailable and only marginal summary moments from published tables of baseline characteristics are available.
Such limitations motivate alternative “balancing” or “calibration” approaches to weighting. These estimate weights under the condition that covariates are balanced, viewing balance as an optimization problem, without explicitly modeling the propensity score. Generally, balancing approaches to weighting are: (1) less susceptible to bias by directly enforcing covariate balance; (2) produce more stable weights, which translate into larger effective sample sizes and more precise treatment effect estimation; and (3) are applicable where only aggregate-level marginal covariate moments are available for the external control.66, 67, 68, 69
Our focus here is on an entropy balancing approach37 called matching-adjusted indirect comparison (MAIC),38, 39, 40 but see Chattopadhyay et al (2020) and Filla et al (2024) for details about similar balancing techniques.66, 67 MAIC is the most popular balancing method in the context of externally controlled SATs. It has many features that are considered desirable: “linear double robustness”, minimally dispersed weights, and the estimation of odds weights that are guaranteed to be positive, resulting in increased interpretability and sample-boundedness, i.e., interpolating the observed data as opposed to extrapolating beyond its support.67, 70 We review the main steps of MAIC, building on prior literature.38, 69, 71, 72, 73, 74
While MAIC does not explicitly model the propensity score as a function of baseline covariates, it implicitly assumes the following logistic model for data source assignment:
$\ln(\omega_{i})\propto\ln\Bigg{(}\frac{(1-e_{i})}{e_{i}}\Bigg{)}=\gamma_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma},$ (8)
where $\omega_{i}$ is a weight proportional to the inverse conditional odds of SAT participation for subject $i=1,\dots,n$, $\gamma_{0}$ is an intercept term parameter and $\mathbf{\gamma}$ is a vector of model parameters. In Equation 8, it is the log-odds of external control participation, $\operatorname{logit}(1-e_{i})$, that are linear across the covariate balance functions. Because $\operatorname{logit}(1-e_{i})=-\operatorname{logit}(e_{i})$, this implies that the log-odds of SAT participation, $\operatorname{logit}(e_{i})$, also vary linearly with $\mathbf{c}(\mathbf{X}_{i})$ as per Equation 2.
Signorovitch et al38 propose using the “method of moments” to estimate the model in Equation 8, such that:
$\frac{\sum_{i=1}^{n_{1}}\omega_{i}\mathbf{c}(\mathbf{X}_{i})}{\sum_{i=1}^{n_{1}}\omega_{i}}$ $=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{c}(\mathbf{X}_{i})$ (9)
[…]
$\frac{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)\mathbf{c}(\mathbf{X}_{i})}{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)}$ $=\hat{\mathbf{\theta}}.$ (11)
where $\hat{\mathbf{\theta}}=(\hat{\mathbf{\theta}}_{1},\hat{\mathbf{\theta}}_{2},\dots,\hat{\mathbf{\theta}}_{k})^{\top}$ is a vector of covariate balance function moments $j=1,\dots,k$ for the external control sample, with $\hat{\mathbf{\theta}}_{j}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}c_{j}(\mathbf{X}_{i})$ assumed to be a consistent estimator for $\mathbf{\theta}_{j}$. Equation 9 is a constraint enforcing that the covariate distributional features of the weighted SAT subjects are exactly balanced with respect to those of the unweighted external control subjects. Equation 10 follows from introducing the assumed model in Equation 8 into the balancing constraint, and Equation 11 results from the exponentiated intercept terms canceling out.
Replacing $\mathbf{\gamma}$ with estimate $\hat{\mathbf{\gamma}}$ in Equation 11 and centering the SAT covariate balance functions on their external control means, one obtains:
$\frac{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)\mathbf{c^{*}}(\mathbf{X}_{i})}{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}=\mathbf{0},$ (12)
where $\mathbf{0}$ is a vector of zeros and $\mathbf{c^{*}}(\mathbf{X}_{i})=\mathbf{c(\mathbf{X}_{i})}-\hat{\mathbf{\theta}}$ is a vector of centered covariate balance functions for subject $i=1,\dots,n_{1}$ in the SAT. Then, because the denominator is positive, Equation 12 is equal to $\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)\mathbf{c^{*}}(\mathbf{X}_{i})=\mathbf{0}$. Solving for $\hat{\mathbf{\gamma}}$ is equivalent to minimizing the objective function:
$Q(\hat{\mathbf{\gamma}})=\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right),$ (13)
as the derivative of $Q(\hat{\mathbf{\gamma}})$ with respect to $\hat{\mathbf{\gamma}}$ is $\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)\mathbf{c^{*}}(\mathbf{X}_{i})$. The objective function in Equation 13 is strictly convex and can be minimized using standard Newton-type convex optimization algorithms,37 yielding an unique finite solution corresponding to the global minimum of $Q(\hat{\mathbf{\gamma}})$.
We have $\omega_{i}\propto\exp\left(\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)\propto\exp\left(\left(\mathbf{c}(\mathbf{X}_{i})-\mathbf{\theta}\right)^{\top}\mathbf{\gamma}\right)=\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\mathbf{\gamma}\right)$. Subject to the normalization constraint $\sum_{i=1}^{n_{1}}\hat{\omega}_{i}=1$, weights for each individual $i=1,\dots,n_{1}$ in the SAT are estimated as:
$\hat{\omega}_{i}=\frac{\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}{\sum_{i=1}^{n_{1}}\exp\left(\mathbf{c^{*}}(\mathbf{X}_{i})^{\top}\hat{\mathbf{\gamma}}\right)}.$ (14)
Note that this definition of the entropy balancing weights coincides with the definitions presented by Jiang et al (2024) 74 and by Jackson et al (2021).73
Similar to Equation 3 and Equation 4, the ATC is estimated by contrasting mean absolute outcomes on the additive scale:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (15)
where the weights have already been normalized to sum to one. Alternatively, fitting a weighted univariable regression of outcome on treatment to the concatenated IPD (with weights $\hat{\omega}_{i}$ for $i=1,\dots,n_{1}$ and unit weights for $i=n_{1}+1,\dots,n$) has been proposed, with the treatment coefficient of the fitted model yielding an ATC estimate.47, 52, 75, 76, 77 We discourage this approach. The weights estimated by Equation 14 are relative; their arbitrary rescaling by a constant of proportionality, e.g., a normalization constant, will also balance the specified covariate functions and does not affect $\hat{\mu}_{0}^{1}$ and $\widehat{\textrm{ATC}}$ in Equation 15.71, 73 Nevertheless, rescaling the weights while retaining weights of one for the external control subjects results in different fitted models.73
The method of moments MAIC estimator gives the Lagrangian dual solution to an entropy balancing primal problem: minimizing the negative entropy of the weights.40, 42, 71 Namely, minimizing the objective function in Equation 13 is equivalent to minimizing the negative entropy $\sum_{i=1}^{n_{1}}\omega_{i}\ln(\omega_{i})$ with Lagrange multipliers.42, 71 The dual optimization problem is easier to solve than the primal37 and has been formulated in different ways.40, 71, 74 These may perform differently computationally – e.g., if minimization is performed on the log scale71 – but result in the same unique dual solution, up to numerical error, due to strict convexity. Because the negative entropy measures the distance of the weights from a uniform distribution, its minimization should produce less disperse weights than the modeling approach in Section 3.2.
While the entropy balancing weights ($\hat{\omega}_{i}$) defined in Equation 14 will be different than the IOW weights ($\widehat{w}_{i}$) obtained from maximum-likelihood estimation of the logistic regression model in Equation 2, $\hat{\nu}_{i}=n_{0}\hat{\omega}_{i}$ will consistently estimate the true inverse odds if the logistic regression model is correctly specified. To be clear, if the logistic regression model correctly specifies the true propensity score model, then we have both: $\widehat{w}_{i}\rightarrow\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}$ (i.e., the IOW weights are consistent) and $\widehat{\nu}_{i}\rightarrow\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}$ (i.e., the normalized entropy balancing weights are consistent), for $i$ in $n_{1}$; see Zhao and Percival (2017) 43 for details. Therefore, following the same logic as detailed in Section 3.2 for the IOW estimators, the entropy balancing estimator, as defined in Equation 15, is also consistent if the implied propensity score model is correctly specified.
In addition, the entropy balancing estimator has a “linear double robustness” property as it is consistent under two distinct underlying data-generating models.40, 42, 43 Namely, it enables consistent estimation as long as either the true log-odds of the propensity score are linear across the specified balance functions (i.e., the implied propensity model is correctly specified); or, the true potential outcome under the active intervention is linear across the specified balance functions. Respectively, either $\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\alpha}$ for some parameters $\alpha_{0}$ and $\mathbf{\alpha}$; or, $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{c}(\mathbf{X}_{i})^{\top}\mathbf{\beta}$ for some parameters $\beta_{0}$ and $\mathbf{\beta}$.
For continuous covariates, it is common practice to only balance first-order moments (means) and, potentially, second-order sample moments (variances).25, 28, 38, 73 Covariate means are balanced by setting $\mathbf{c}(\mathbf{X}_{i})=c_{1}(\mathbf{X}_{i})=\mathbf{X}_{i}$ and $\hat{\mathbf{\theta}}=\hat{\mathbf{\theta}}_{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}$, in which case consistency is guaranteed as long as either the true propensity score model is $\operatorname{logit}(e_{i})=\alpha_{0}+\mathbf{X}_{i}^{\top}\mathbf{\alpha}$ for some parameters $\alpha_{0}$ and $\mathbf{\alpha}$; or, the potential outcome for the active intervention varies linearly with the covariates $\mathbf{X}_{i}$ that are mean-balanced: $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{X}_{i}^{\top}\mathbf{\beta}$ for some parameters $\beta_{0}$ and $\mathbf{\beta}$.
Both covariate means and variances are balanced by setting $\mathbf{c}(\mathbf{X}_{i})=[c_{1}(\mathbf{X}_{i}),c_{2}(\mathbf{X}_{i})]^{\top}=[\mathbf{X}_{i},\mathbf{X}^{2}_{i}]^{\top}$, $\hat{\mathbf{\theta}}=(\hat{\mathbf{\theta}}_{1},\hat{\mathbf{\theta}}_{2})^{\top}$, $\hat{\mathbf{\theta}}_{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}$ and $\hat{\mathbf{\theta}}_{2}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\mathbf{X}_{i}^{2}$, in which case consistency is achieved if either the true propensity score model is $\textrm{logit}(e_{i})=\alpha_{0}+\mathbf{X}_{i}^{\top}\mathbf{\alpha_{1}}+(\mathbf{X}_{i}^{2})^{\top}\mathbf{\alpha_{2}}$ for some parameters $\alpha_{0}$, $\mathbf{\alpha_{1}}$ and $\mathbf{\alpha_{2}}$; or, the potential outcome for the active intervention varies quadratically with $\mathbf{X}_{i}$, such that $\textrm{E}(Y^{1}_{i}\mid\mathbf{X}_{i})=\beta_{0}+\mathbf{X}_{i}^{\top}\mathbf{\beta}_{1}+(\mathbf{X}_{i}^{2})^{\top}\mathbf{\beta}_{2}$ for some parameters $\beta_{0}$, $\mathbf{\beta_{1}}$ and $\mathbf{\beta_{2}}$. Balancing the means of the covariates and the squared covariates enforces that variances are balanced because $\textrm{Var}(X)=\textrm{E}(X^{2})-\textrm{E}(X)^{2}$.
3.4 Doubly robust augmented weighting estimators
Unfortunately, the aforementioned balancing strategies are insufficient for consistency where the underlying data-generating models are more complex. One can consider various other non-linear transformations of $\mathbf{X}_{i}$, e.g., higher-order polynomial terms and flexible basis functions such as splines, and balance on the means of the transformed covariates. This would enable one to conjecture more flexible implicit data-generating models that are non-linear in the covariates. Moreover, one could go beyond the marginal moments of individual covariates and also balance joint covariate distributions; for instance, by including quantiles of interactions for pairs or triples of covariates.78 However, pursuing these more ambitious balancing strategies is often infeasible:
  • As the number of balancing conditions increases, it is more likely that $\mathbf{\theta}$ falls outside the convex hull of $\mathbf{c}(\mathbf{X}_{i})$ for $i=1,\dots,n_{1}$.60 This implies that a feasible weighting solution to the convex optimization problem does not exist: there is no set of positive weights that can enforce balance in the required distributional features and the numerical optimization algorithm will fail to converge.66
  • Increasing the number of balancing conditions leads to further reductions in effective sample size and precision, which are particularly problematic with low sample sizes and poor covariate overlap.
  • Where covariate IPD for the external control are unavailable, only first- and second-order marginal moments from published tables of baseline characteristics are often available for balancing. Higher-order moments and the means of transformed covariates are rarely reported.
These difficulties motivate the explicit augmentation of the weighting estimators, as proposed by Funk et al (2011),79 allowing for a less restrictive outcome model that permits non-linear link functions and/or outcomes that depend on non-linear covariate transformations.
To construct such augmented estimators, we first postulate a model $m\left(\mathbf{X}_{i};\mathbf{\beta}\right)$ for the potential outcome expectation under the active intervention, conditional on covariates $\mathbf{X}_{i}$:
$q\left(\textrm{E}(Y_{i}^{1}\mid\mathbf{X}_{i};\mathbf{\beta})\right)=m\left(\mathbf{X}_{i};\mathbf{\beta}\right),$ (16)
where $q(\cdot)$ denotes an appropriate link function and $\mathbf{\beta}$ is a vector of model parameters encoding the covariate-outcome relationships. For instance, a logistic regression could be used for binary outcomes, such that the model is binomial, the link function is the logit and the potential outcome expectation is on the probability scale. We have assumed that the model for the conditional outcome expectation is parametric, but this need not necessarily be the case.80
The model in Equation 16 is fitted to the SAT participants $i=1,\dots,n_{1}$ using maximum-likelihood estimation, such that the fitted model $m\left(\mathbf{X}_{i};\hat{\mathbf{\beta}}\right)$ with parameter estimates $\hat{\mathbf{\beta}}$ is an estimator of the (transformed) conditional expectation $q\left(\textrm{E}(Y_{i}^{1}\mid\mathbf{X}_{i},S_{i}=1)\right)$. Based on the model, potential outcomes under the active intervention are predicted for each subject in the external control:
$\hat{Y}_{i}^{1}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right).$ (17)
In this case, $i=n_{1}+1,\dots,n$, and the predicted outcomes are counterfactual because the subjects in the external control have not received the active intervention.
By averaging or “marginalizing” the potential outcome predictions generated by Equation 17 over the empirical covariate distribution of the external control arm, one obtains the “G-computation”, “regression standardization” or “plug-in G-formula” estimator33, 34, 44, 45, 46 for the mean potential outcome $\mu_{0}^{1}$ had subjects in the external control received the active intervention:
$\hat{\mu}_{0}^{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y}_{i}^{1}.$ (18)
The G-computation estimator for the ATC contrasts the average of potential counterfactual outcomes under the active intervention with the average of observed outcomes for the external control. We convert mean absolute outcomes to the additive scale imposed by link function $g(\cdot)$, prior to taking the difference between treatments on such scale:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y}_{i}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}}.$ (19)
Note that the link function $g(\cdot)$ used for summarizing the treatment effect does not necessarily need to match the link function $q(\cdot)$ used for modeling.81
The G-computation estimators in Equations 18 and 19 rely on the outcome model in Equation 16 being correctly specified, in order to be consistent for the mean absolute outcome ${\mu}_{0}^{1}$ and the ATC, respectively. Assuming correct model specification for all estimators, G-computation is more precise and efficient than weighting, particularly when poor overlap leads to large reductions in effective sample size.34, 46 However, the increase in precision is achieved by implicit extrapolation into non-overlapping regions of the covariate space, hiding underlying failures of the positivity assumption.48 Model misspecification bias is almost impossible to diagnose in extrapolated regions, and there is typically no inflation of the variance to reflect the extrapolation uncertainty. Our proposed augmented weighting estimators will not use the outcome model to extrapolate, but to gain bias-robustness – and, potentially, some precision35 – with respect to their weighting counterparts.
Firstly, based on Funk et al (2011) and Shinozaki and Matsuyama (2015),79, 82 we propose combining the modeling approach to inverse odds weighting, described in Section 3.2, with G-computation. Specifically, suppose we have fitted the outcome model in Equation 16 to the SAT. We now let $q^{-1}\left(m(\mathbf{X}_{i};\hat{\mathbf{\beta}})\right)$ be a prediction of potential outcome $Y_{i}^{1}$ for the active intervention based on the fitted regression, not only for the external control subjects but for all subjects $i=1,\dots,n$ in the SAT and the external control. The propensity score weights $\hat{w}_{i}$ derived for $i=1,\dots,n_{1}$ are used to add an error-correcting term to the G-computation estimator. The resulting augmented inverse odds weighting estimator for the mean absolute outcome $\mu_{0}^{1}$ is:
$\hat{\mu}_{0}^{1}=$ $\frac{1}{n_{0}}\left(\sum_{i=1}^{n_{1}}\hat{w}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}$
$=$ $\frac{1}{n_{0}}\left(\sum_{i=1}^{n_{1}}\hat{w}_{i}\epsilon_{i}^{1}+\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right),$ (20)
where $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ is a residual term for subject $i=1,\dots,n_{1}$ in the SAT. Note that this estimator exactly corresponds to the doubly robust estimator proposed by Shinozaki and Matsuyama (2015),82 except that their estimator is for $\textrm{E}(Y^{0}|S=1)$ while the estimator in Equation 20 is for $\textrm{E}(Y^{1}|S=0)$ (the estimand ultimately targeted by Shinozaki and Matsuyama is the ATT but ours is the ATC).
In Equation 20, the G-computation estimator has been augmented with a weighted average of the residuals for the SAT subjects. Such a weighted average is a “one-step” correction term for the potential bias of the G-computation estimator.83 The corresponding bias-corrected estimator for the ATC is:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{1}{n_{0}}\left(\sum_{i=1}^{n_{1}}\hat{w}_{i}\epsilon_{i}^{1}+\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (21)
on the additive scale imposed by link function $g(\cdot)$. Normalizing the weights so that they sum to one and to ensure bounded estimates, we obtain the alternative estimators:
$\hat{\mu}_{0}^{1}=\frac{\sum_{i=1}^{n_{1}}\hat{w}_{i}\epsilon_{i}^{1}}{\sum_{i=1}^{n_{1}}\hat{w}_{i}}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1},$ (22)
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{\sum_{i=1}^{n_{1}}\hat{w}_{i}\epsilon_{i}^{1}}{\sum_{i=1}^{n_{1}}\hat{w}_{i}}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (23)
Equations 22 and 23 should provide improved finite sample properties and more stable and precise estimation than Equations 20 and 21, respectively.
Our novel contribution is combining the entropy balancing-based MAIC approach, described in Section 3.3, with the G-computation estimator. Again, based on the outcome model fitted to the SAT, let $q^{-1}\left(m(\mathbf{X}_{i};\hat{\mathbf{\beta}})\right)$ be a prediction of the potential outcome $Y_{i}^{1}$ under the active intervention for all subjects $i=1,\dots,n$ in the SAT and the external control. In this case, it is the MAIC weights, $\hat{\omega}_{i}$ for $i=1,\dots,n_{1}$, derived in Equation 14, that are used to add an error-correcting term to the G-computation estimator. The resulting augmented MAIC estimator of the mean absolute outcome $\mu_{0}^{1}$ is:
$\hat{\mu}_{0}^{1}=$ $\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}$
$=$ $\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\epsilon_{i}^{1}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1},$ (24)
where the G-computation estimator has been augmented with a weighted average of the residuals $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ for $i=1,\dots,n_{1}$, but this time the weighted average has been computed using the MAIC weights.
The corresponding bias-corrected augmented MAIC estimator for the ATC is:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\epsilon_{i}^{1}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}},$ (25)
on the additive scale imposed by link function $g(\cdot)$. We conjecture that the augmented MAIC estimators in Equations 24 and 25 will perform better statistically than the augmented estimators based on the modeling approach to weighting, which could exhibit erratic performance with highly variable weights, particularly if these are combined with a misspecified outcome model.84, 85 We expect the augmented MAIC estimators to inherit the more attractive properties of the entropy balancing weights: (1) lower susceptibility to bias by directly enforcing covariate balance; and (2) greater stability, translating into larger effective sample sizes after weighting and enhanced precision in estimation.
The augmented weighting estimators in Equations 20, 22 and 24 are doubly robust for the mean potential outcome $\mu^{1}_{0}$. That is, they estimate $\mu^{1}_{0}$ consistently as long as either the propensity score model for data source assignment or the outcome model is correctly specified, but not necessarily both. By extension, the augmented weighting estimators in Equations 21, 23 and 25 are doubly robust for the ATC, under the assumption that $\hat{\mu}_{0}^{0}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}$ is consistent for $\mu_{0}^{0}$. We provide an intuitive heuristic to demonstrate double robustness.
Consider the simple scenario where we have a binary outcome, $Y$, and a single discrete covariate, $X$, and the augmented weighting estimator in Equation 24 is re-written as:
$\hat{\mu}_{0}^{1}=\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\left(Y_{i}-\hat{Y}^{1}_{i}\right)+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}.$ (26)
If the outcome model is correctly specified, the expectation of the first summation in Equation 26 converges to zero as $n_{1}\rightarrow\infty$ because $\hat{Y}_{i}^{1}\rightarrow Y_{i}$ and the terms inside the summation cancel out, irrespective of any postulated propensity score model. The second summation is equivalent to the G-computation estimator and is consistent for $\mu_{0}^{1}$ because the outcome model is correct. Consequently, $\hat{\mu}_{0}^{1}\rightarrow\mu_{0}^{1}$, and $\rightarrow\textrm{ATC}$ ATC (assuming $\hat{\mu}_{0}^{0}\rightarrow\mu_{0}^{0}$).
Conversely, if the propensity score model is correctly specified but the outcome model is incorrect, the first summation consistently cancels out the bias produced by the G-computation estimator in the second summation and the remainder term is exactly equal to the non-augmented weighting estimator, which converges to $\mu_{0}^{1}$ as $n_{1}\rightarrow\infty$ because the propensity score model is correct. To illustrate this, consider rearranging Equation 26 to:
$\hat{\mu}_{0}^{1}$ $=\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}-\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\hat{Y}^{1}_{i}+\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}$
$=\underbrace{\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}}_{C}+\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y_{i}}^{1}-\sum_{i=1}^{n_{1}}\hat{\omega}_{i}\hat{Y}^{1}_{i}\right)}_{D}.$ (27)
First, $C$ is equivalent to the MAIC estimator and is consistent for $\mu_{0}^{1}$ because the propensity score model is correct. Expanding the summations in $D$ over $i=1,...,n$, we have:
$D$ $=\left(\sum_{i=1}^{n}\frac{(1-S_{i})\hat{Y}_{i}^{1}}{n_{0}}-\frac{S_{i}\hat{\nu}_{i}\hat{Y}_{i}^{1}}{n_{0}}\right)$ (28)
$=\frac{1}{n_{0}}\left(\sum_{i=1}^{n}{\hat{Y}_{i}^{1}}\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\right),$ (29)
where $\hat{\nu}_{i}=n_{0}\hat{\omega}_{i}$, for $i$ in 1,…,$n$. Then, suppose that the outcome model for $Y^{1}$ is independent of $S$ and that $Y^{1*}_{i}$ is the large sample limit of $\hat{Y}^{1}_{i}$. Also, suppose that the large sample limit of $\hat{\nu}_{i}$ is $\frac{\textrm{Pr}(S=0|X=x_{i})}{\textrm{Pr}(S=1|X=x_{i})}$, since the propensity score model is correctly specified. Then:
$\textrm{E}\left(\hat{Y}_{i}^{1}\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\right)$ $=\textrm{E}\left(\textrm{E}\left(\hat{Y}_{i}^{1}\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\mid X\right)\right)$
$=\textrm{E}\left(\textrm{E}(\hat{Y}_{i}^{1}\mid X)\times\textrm{E}\left(\left((1-S_{i})-S_{i}\hat{\nu}_{i}\right)\mid X\right)\right)$
$\rightarrow\textrm{E}\left(Y^{1*}_{i}\sum_{x}{\textrm{Pr}(S=0|X=x)\textrm{Pr}(X=x)-\textrm{Pr}(S=1|X=x)\textrm{Pr}(X=x)\frac{\textrm{Pr}(S=0|X=x)}{\textrm{Pr}(S=1|X=x)}}\right)$
$=\textrm{E}\left(Y^{1*}_{i}\sum_{x}{\textrm{Pr}(S=0|X=x)\textrm{Pr}(X=x)-\textrm{Pr}(X=x)\textrm{Pr}(S=0|X=x)}\right)$
$=\textrm{E}\left(Y^{1*}_{i}\times 0\right)$
$=0.$ (30)
Consequently, $D\rightarrow 0$ which implies that $\hat{\mu}_{0}^{1}\rightarrow\mu_{0}^{1}$, and $\rightarrow\textrm{ATC}$ ATC having assumed $\hat{\mu}_{0}^{0}\rightarrow\mu_{0}^{0}$.
While all the augmented weighting estimators described in this section are doubly robust, the augmented MAIC estimator defined in Equation 25 is arguably more robust to model misspecification bias because it is consistent under a greater number of distinct underlying data-generating mechanisms. Namely, the augmented MAIC estimator is consistent as long as either: (1) the log-odds of the propensity score are linear across the covariate balance functions; (2) the potential outcome under the active intervention is linear across the covariate balance functions; or (3) the explicit augmentation model for the potential outcome under the active intervention is correctly specified. The estimation of the weights is consistent as long as either the first or the second condition holds. Conversely, the augmented estimators based on the modeling approach to weighting (defined in Equations 20 to 23) are consistent as long as either the first or the third condition holds, with the first condition being necessary for consistent estimation of the weights.
3.5 Other augmented weighting estimators
In Section 3.4, we proposed augmented estimators that combine the predictions of an unweighted outcome model with weights in a weighted average. Nevertheless, there are other ways of constructing augmented weighting estimators. One approach claimed to be doubly robust consists of G-computation based on the predictions of a weighted outcome model.52, 61 Where the target estimand is the ATC, this involves: (1) estimating weights using the methods described in Section 3.2 and Section 3.3; (2) fitting a weighted model for the conditional outcome expectation to the SAT participants; and (3) marginalizing the outcome predictions of the weighted regression over the external control covariate distribution. The resulting estimator for the mean absolute outcome $\mu_{0}^{1}$ would be:
$\hat{\mu}_{0}^{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\hat{Y}_{i}^{1}=\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right),$ (31)
where $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})$ indexes the fitted weighted regression with vector $\hat{\mathbf{\beta}}_{v}$ of parameter estimates. The ATC would be estimated by substituting Equation 31 into Equation 19.
Such estimators are only doubly robust where the outcome model is a generalized linear model (GLM) with a canonical link function $q(\cdot)$.51, 61, 84, 86 The estimator in Equation 31 and the corresponding ATC estimator would not be doubly robust where the GLM link function is non-canonical,51 or where the outcome model is a Cox proportional hazards model or a parametric survival model in the time-to-event setting.87 Nevertheless, results by Gabriel et al (2024) suggest asymptotic equivalence and similar finite-sample performance to the augmented weighting estimators in Section 3.4 for GLMs with canonical link functions fitted via maximum-likelihood,51 provided that the same weights are used and correct model specification. We note that the target of the investigations by Gabriel et al (2024) is the ATE and the modeling approach to weighting.51
3.6 Variance estimation
To estimate the variance and construct confidence intervals for $\hat{\mu}_{0}^{1}$ and $\widehat{\textrm{ATC}}$, it is possible to use empirical sandwich-type (“robust”) variance estimators to account for the correlation induced by weighting.25, 38 In the specific context of non-randomized comparisons, such as the externally controlled SATs and unanchored ITCs explored in this article, these estimators have exhibited either under-precision or over-precision for the ATT (or the ATC)88, 89 and under-precision for the ATE.36, 65 This is because most implementations ignore the estimation of the propensity score model or the weights, assuming the weights to be fixed quantities.88, 89
Analytic expressions that incorporate weight estimation could be derived,39, 88, 89 but we propose a practical alternative based on the ordinary non-parametric bootstrap,36 explicitly accounting for uncertainty in the weight estimation. This involves resampling with replacement the concatenated IPD consisting of the SAT and external control data. In each bootstrap iteration, the weight estimation and/or outcome modeling procedures are performed, and $\mu_{0}^{1}$, $\mu_{0}^{0}$ and ATC are re-estimated. Standard errors for $\hat{\mu}_{0}^{1}$, $\hat{\mu}_{0}^{0}$ and $\widehat{\textrm{ATC}}$, are given by the standard deviations across the bootstrap resamples. Subsequently, Wald-type confidence intervals can be constructed. Alternatively, one can directly calculate confidence intervals from the percentiles of the bootstrap resamples, e.g., 2.5% and 97.5% for the 95% confidence interval.
3.7 External controls with unavailable individual participant data
In the context of unanchored ITCs, the external control is often a historical trial for which IPD are unavailable, due to privacy or confidentiality reasons. In this case, only published AD are available for the external control.25, 28, 32 Such data consist of marginal summary moments $\hat{\mathbf{\theta}}$ from reported tables of baseline characteristics, typically only including means and standard deviations (for continuous covariates), and an estimate $\hat{\mu}_{0}^{0}$ of the mean absolute outcome under the control in the external data source.25, 28, 38, 69, 73 An important shortfall of this scenario is the need to assume that $\hat{\mathbf{\theta}}_{j}=\mathbf{\theta}_{j}$ with zero variability for the covariate balance function moments $j=1,\dots,k$, i.e., that the external control covariate distributional data are fixed.40 While this may be reasonable with large sample sizes for the external control, it can otherwise result in overly precise inferences and inflated Type I error rates.40
In this setting, for all methods except (non-augmented) MAIC, one must simulate $M$ individual-level covariate profiles from the assumed covariate distribution of the external control based on published summary statistics.46, 81, 90 The number $M$ of hypothetical subject profiles should be relatively large, e.g., $M=1000$, to minimize sampling variability and random seed sensitivity, and does not necessarily need to match the original sample size $n_{0}$ of the external control.46, 81, 90 Necessary information to infer the joint covariate distribution of the external control, e.g., distributional forms and correlation structures, is rarely published. Hence, this must be borrowed from other data sources or selected based on theoretical properties, following recommendations in the literature.25, 28, 46, 81, 90
The notation and procedures for Section 3.2, Section 3.4 and Section 3.5 change as follows. The observed IPD for the SAT is stacked with the simulated subject-level covariate data for the external control. The concatenated dataset is now $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$ for $i=1,\dots,n_{1},n_{1}+1,\dots,n_{1}+M$. For the SAT subjects $i=1,\dots,n_{1}$, we have $S_{i}=1$ and $T_{i}=1$, with $\mathbf{X}_{i}$ and $Y_{i}$ corresponding to the actual covariate and outcome values observed in the trial. For the hypothetical external controls $i=n_{1}+1,\dots,n_{1}+M$, we have $S_{i}=0$ and $T_{i}=0$, the values of $\mathbf{X}_{i}$ are simulated, and $Y_{i}$ are unavailable (but not required for the analysis given that the target estimand is the ATC and that estimate $\hat{\mu}_{0}^{0}$ is available from published results).
The general form of the ATC inverse odds weighting estimators described in Section 3.2 is now:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{1}{K}\sum_{i=1}^{n_{1}}\hat{w}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right),$
where $K$ is a constant. There is only a change in notation here given that individual-level outcomes under the control are now unavailable for the external data source, and cannot be included in the concatenated dataset.
For the G-computation ATC estimator outlined in Section 3.4, we now have:
$\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y}_{i}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right),$ (32)
where the potential outcome predictions $\hat{Y}_{i}^{1}$ under the active intervention are generated for each hypothetical external control subject $i=n_{1}+1,\dots,n_{1}+M$, and averaged over the simulated covariate profiles.
The general form of our doubly robust augmented weighting ATC estimators, proposed in Section 3.4, is now:
[…]
where $K$ is a constant, $\epsilon_{i}^{1}=Y_{i}-\hat{Y}_{i}^{1}$ and $\hat{v}_{i}$ are a residual term and a weight estimate, respectively, for $i=1,\dots,n_{1}$, and the potential outcome predictions $\hat{Y}_{i}^{1}$ under the active intervention are generated for all SAT subjects and hypothetical external controls $i=1,\dots,n_{1},n_{1}+1,\dots,n_{1}+M$.
For the augmented “weighted G-computation” estimator in Section 3.5, the outcome predictions would be averaged over the simulated covariates for the external control. The resulting estimator for the mean absolute outcome $\mu_{0}^{1}$ is $\hat{\mu}_{0}^{1}=\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y}_{i}^{1}=\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right)$, which would be input into Equation 32 for estimation of the ATC.
The unavailability of IPD for the external control entails some changes to the non-parametric bootstrap procedure described in Section 3.6 for variance estimation. In this case, only the SAT data, $(S_{i},\mathbf{X}_{i},T_{i},Y_{i})$ for $i=1,\dots,n_{1}$, are resampled to re-estimate $g(\mu_{0}^{1})$ in each bootstrap iteration, with the standard error, $\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)$, computed as the standard deviation over the bootstrap resamples. Then, the decomposition:
SE(ATC^)=($\textrm{SE}\left(\widehat{\textrm{ATC}}\right)=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}},$ (33)
is used to estimate the standard error of the ATC, where $\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)$ is derived from published aggregate-level results.46, 77, 90 A limitation of the above formula is that it assumes that the mean absolute outcomes are statistically independent. Moreover, while computing $\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)$ is a trivial exercise for continuous and binary outcomes, (e.g., there is a closed-form formula for the standard error of the log-odds using the Delta method), it can be challenging for other outcomes such as those in the time-to-event setting.77, 90 Once $\textrm{SE}\left(\widehat{\textrm{ATC}}\right)$ is computed, Wald-type confidence intervals can be readily constructed.
3.8 Targeting the average treatment effect in the treated
We briefly adapt the methodologies in Section 3.2 to Section 3.5 so that these target the ATT. We assume that there is full IPD availability and that $\hat{\mu}_{1}^{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}$ is consistent for $\mu_{1}^{1}$.
For the modeling-based approach to odds weighting in Section 3.2, external control subjects are weighted by their conditional odds of SAT participation to transport the external control outcomes to the SAT (sub) population. SAT subjects are unweighted and external control subjects $i=n_{1}+1,\dots,n$ are weighted by $\hat{w}_{i}=\hat{e}_{i}/(1-\hat{e}_{i})$. Assuming correct specification of the propensity score model, the estimated weights would balance the covariate distribution of the external control with respect to that of the SAT, enabling consistent estimation of mean absolute outcome $\mu^{0}_{1}$ and the ATT. Propensity score predictions that are close to one lead to extreme weights and imprecise ATT estimation, particularly where the sample size of the external control is small.
A MAIC estimator for the ATT, akin to that described in Section 3.3, would enforce that the covariate distributional features of the weighted external control subjects are exactly balanced with respect to those of the SAT subjects. As such, the balancing constraints would center the external control covariate balance functions on their SAT means. MAIC enables consistent estimation of $\mu^{0}_{1}$ and the ATT, as long as either the log-odds of the propensity score or the potential outcome under the control are linear across the specified balance functions. The general form of the weighting estimators for the ATT is:
$\widehat{\textrm{ATT}}=g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}\right)}_{\hat{\mu}_{1}^{1}}-g\underbrace{\left(\frac{1}{K}\sum_{i=n_{1}+1}^{n}\hat{v}_{i}Y_{i}\right)}_{\hat{\mu}_{1}^{0}},$
where $K$ is a constant and $\hat{v}_{i}$ is a weight estimate for $i=n+1,\dots,n$, derived using the modeling approach or MAIC.
A G-computation estimator such like that described in Section 3.4 but for the ATT requires postulating a model for the potential outcome expectation under the control, fitted to the external control participants. Based on the fitted model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})$, the potential outcome under the control is predicted for each subject $i=1,\dots,n_{1}$ in the SAT: $\hat{Y}_{i}^{0}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$. The potential outcome predictions are averaged over the empirical covariate distribution of the SAT, resulting in the ATT estimator:
$\widehat{\textrm{ATT}}=g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}\right)}_{\hat{\mu}_{1}^{1}}-g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y}^{0}_{i}\right)}_{\hat{\mu}_{1}^{0}},$ (34)
which relies on correct specification of the model for the potential outcome under the control for consistent estimation.
Our doubly robust augmented weighting estimators, proposed in Section 3.4, would target the ATT as follows. Based on an outcome model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})$ fitted to the external control participants, the potential outcome under the control is predicted for all subjects $i=1,\dots,n$·in the SAT and the external control: $\hat{Y}_{i}^{0}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$. The potential outcome predictions are augmented with a weighted average of residuals for the external control subjects. The general form of the doubly robust augmented weighting estimators for the ATT is:
$\widehat{\textrm{ATT}}=g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}\right)}_{\hat{\mu}_{1}^{1}}-g\underbrace{\left(\frac{1}{K}\sum_{i=n_{1}+1}^{n}\hat{v}_{i}\epsilon_{i}^{0}+\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y_{i}}^{0}\right)}_{\hat{\mu}_{1}^{0}},$
where $K$ is a constant, $\hat{v}_{i}$ is a weight estimate and $\epsilon_{i}^{0}=Y_{i}-\hat{Y}_{i}^{0}$ is a residual term for subject $i=n_{1}+1,\dots,n$ in the external control.
An augmented “weighted G-computation” estimator akin to that described in Section 3.5 would target the ATT by: (1) estimating weights for the odds of SAT participation; (2) fitting a weighted model $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})$ for the conditional outcome expectation to the external control participants; and (3) averaging the outcome predictions of the weighted regression over the SAT covariate distribution. The resulting estimator for the mean absolute outcome $\mu_{1}^{0}$ is $\hat{\mu}_{1}^{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\hat{Y}_{i}^{0}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}}_{v})\right)$, which is then substituted into Equation 34 for estimation of the ATT.
4 Simulation study
4.1 Aims
We conducted a simulation study to evaluate the performance of various estimators under different conditions. The simulation study design was planned following the structured “ADEMP” approach outlined by Morris et al (2019),91 to ensure reproducibility and meaningful conclusions. Specifically, we clearly defined research aims, data-generating mechanisms under controlled scenarios and estimands, and assessed the performance of several estimators using relevant performance measures: bias, variance and coverage. All simulations and analyses were performed using R statistical software version 4.3.1.92 The files and code required to run the simulations are publicly available on Github at https://github.com/harlanhappydog/DRAWE-.
4.2 Data-generating mechanisms
We simulated data inspired by the data-generating mechanisms in a simulation study by Kang and Schafer (2007).85 Some modifications were required since Kang and Schafer (2007) considered continuous-valued outcomes,85 while we consider binary outcomes. The simulated data consist of $\{\mathbf{X}_{i},\mathbf{Z}_{i},T_{i},S_{i},Y_{i}\},i=1,\ldots,n\}$, with the control group fully external such that $S_{i}=T_{i}$, with $n_{1}=\sum_{i=1}^{n}{S_{i}}$, and $n_{0}=\sum_{i=1}^{n}{(1-S_{i})}$, as detailed in Section 3.1. While $\mathbf{X}_{i}$ is observed, $\mathbf{Z}_{i}$ is unobserved. To generate the data, $\mathbf{X}_{i}$ is distributed as $\mathrm{Normal}(0,I_{4})$, for $i$ in 1,…,$n$, and $\mathbf{Z}_{i}$ is then obtained by applying the following transformations:
$Z_{i1}$ $=\textrm{scale}(\exp(X_{i1}/2)),$
$Z_{i2}$ $=\textrm{scale}(X_{i2}^{2}),$
$Z_{i3}$ $=\textrm{scale}((X_{i1}X_{i3}+0.6)^{3}),$
$Z_{i4}$ $=\textrm{scale}((X_{i2}+X_{i4}+20)^{2}).$
where $\textrm{scale}(\dot{)}$ indicates normalization such that $Z_{i1},Z_{i2},Z_{i3}$ and $Z_{i4}$ each have mean of 00 and standard deviation of $1$. Note that these transformations for deriving $\mathbf{Z}_{i}$ from $\mathbf{X}_{i}$ are similar to the ones detailed by Kang and Schafer (2007)85 but not identical, with changes made to highlight the consequences of model misspecification.
We consider four different scenarios. For each, we generated 10,000 simulated datasets of size $n=200$ and $n=1000$. Note that in all four scenarios the distribution of $S$ is approximately balanced such that $n_{1}\approx n_{0}$. The four scenarios are defined as:
  • KS1: $Y_{i}$ is generated from a Bernoulli distribution with
    $\quad\textrm{Pr}(Y_{i}=1\mid\mathbf{X}_{i},T_{i})=\mathrm{expit}({X}_{1i}-1.50{X}_{2i}+0.5{X}_{3i}-0.5{X}_{4i}+1.50T_{i}-0.50T_{i}{X}_{1i})$,
    where $T_{i}=S_{i}$, and $S_{i}$ is generated from a Bernoulli distribution with
    $\textrm{Pr}(S_{i}=1\mid\mathbf{X}_{i})=\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.5X_{i4})$.
    In this scenario, both the outcome model and the propensity score model for data source assignment are correctly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.68, 0.85, 0.92, and 0.85 for $X_{1}$,$X_{2}$,$X_{3}$,and $X_{4}$, respectively, see Figure 3 in the Supplementary Material.
  • KS2: $Y_{i}$ is generated from a Bernoulli distribution with
    $\quad\textrm{Pr}(Y_{i}=1\mid\mathbf{Z}_{i},T_{i})=\mathrm{expit}({Z}_{1i}-1.50{Z}_{2i}+0.5{Z}_{3i}-0.5{Z}_{4i}+1.50T_{i}-0.50T_{i}{Z}_{1i})$,
    where $T_{i}=S_{i}$, and $S_{i}$ is generated from a Bernoulli distribution with
    $\textrm{Pr}(S_{i}=1\mid\mathbf{X}_{i})=\mathrm{expit}(-X_{i1}+0.5X_{i2}-0.25X_{i3}-0.5X_{i4})$.
    The relevant covariate adjustment approaches would fit an outcome model (of the same form) to the observed $\mathbf{X}_{i}$, as the $\mathbf{Z}_{i}$ used for the true outcome-generating process are unobserved. This implies that the outcome model is incorrectly specified. The propensity score model is still correctly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.68, 0.85, 0.92, and 0.85 for $X_{1}$,$X_{2}$,$X_{3}$,and $X_{4}$, respectively, see Figure 3 in the Supplementary Material.
  • KS3: $Y_{i}$ is generated from a Bernoulli distribution with
    $\quad\textrm{Pr}(Y_{i}=1\mid\mathbf{X}_{i},T_{i})=\mathrm{expit}({X}_{1i}-1.50{X}_{2i}+0.5{X}_{3i}-0.5{X}_{4i}+1.50T_{i}-0.50T_{i}{X}_{1i})$,
    where $T_{i}=S_{i}$, and $S_{i}$ is generated from a Bernoulli distribution with
    $\textrm{Pr}(S_{i}=1\mid\mathbf{Z}_{i})=\mathrm{expit}(-Z_{i1}+0.5Z_{i2}-0.25Z_{i3}-0.5Z_{i4})$.
    The relevant covariate adjustment approaches would fit a propensity score model (of the same form) to the observed $\mathbf{X}_{i}$, as the $\mathbf{Z}_{i}$ used for the true data source assignment process are unobserved. This implies that the propensity score model is incorrectly specified. The outcome model is correctly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.71, 0.84, 0.99, and 0.89 for $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively, see Figure 4 in the Supplementary Material.
  • KS4: $Y_{i}$ is generated from a Bernoulli distribution with
    $\quad\textrm{Pr}(Y_{i}=1\mid\mathbf{Z}_{i},T_{i})=\mathrm{expit}({Z}_{1i}-1.50{Z}_{2i}+0.5{Z}_{3i}-0.5{Z}_{4i}+1.50T_{i}-0.50T_{i}{Z}_{1i})$,
    where $T_{i}=S_{i}$, and $S_{i}$ is generated from a Bernoulli distribution with
    $\textrm{Pr}(S_{i}=1\mid\mathbf{Z}_{i})=\mathrm{expit}(-Z_{i1}+0.5Z_{i2}-0.25Z_{i3}-0.5Z_{i4})$.
    This scenario implies that both the outcome model and the propensity score model are incorrectly specified. The distribution of the covariates is such that overlap across the two groups is relatively high, with overlap proportions of 0.71, 0.84, 0.99, and 0.89 for $X_{1}$, $X_{2}$, $X_{3}$, and $X_{4}$, respectively, see Figure 4 in the Supplementary Material.
4.3 Estimands
The estimand of interest is the ATC, as defined in Section 2. We adopt the logit link function $g(p)=\ln(p/(1-p))$ for marginal outcome probability $p$, such that the ATC is on the marginal log-odds ratio scale. The values of the ATC estimands were calculated numerically, by simulating 1 million binary outcomes using the true data-generating mechanisms outlined in Section 4.2. Data-generating mechanisms KS1, KS2, KS3 and KS4 correspond to true ATCs of 1.128, 1.214, 1.053 and 1.163, respectively.
4.4 Methods
We compared the ten following estimators:
  1. 1.
    The naïve estimator, which does not perform covariate adjustment:
    $\widehat{\textrm{ATC}}=g\underbrace{\left(\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}Y_{i}\right)}_{\hat{\mu}_{1}^{1}}-g\underbrace{\left(\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}Y_{i}\right)}_{\hat{\mu}_{0}^{0}}.$ (35)
  2. 2.
    The IOW estimator with weights derived using the “modeling” approach, as per Equation 3 (Section 3.2).
  3. 3.
    The IOW estimator with normalized weights derived using the “modeling” approach, as per Equation 4 (Section 3.2).
  4. 4.
    The MAIC (entropy balancing) estimator, as per Equation 15 (Section 3.3).
  5. 5.
    The G-computation estimator, as per Equation 19 (Section 3.4).
  6. 6.
    The doubly robust (DR) augmented weighting estimator with “modeling” IOW weights, as per Equation 21 (see Section 3.4).
  7. 7.
    The doubly robust (DR) augmented weighting estimator with normalized “modeling” IOW weights, as per Equation 23 (see Section 3.4).
  8. 8.
    The doubly robust (DR) augmented weighting estimator with MAIC (entropy balancing) weights, as per Equation 25 (our main contribution, see Section 3.4).
  9. 9.
    The augmented “weighted G-computation” estimator described in Section 3.5, said to be doubly robust for the ATE where the outcome model is a GLM with a canonical link function, using the (normalized) “modeling” IOW weights.
  10. 10.
    The augmented “weighted G-computation” estimator described in Section 3.5, but using the MAIC (entropy balancing) weights instead of the “modeling” IOW weights.
For all estimators, we computed 95% confidence intervals using the non-parametric bootstrap approach described in Section 3.6. Specifically, we used $B=100$ bootstrap resamples of the concatenated SAT and external control IPD to approximate the standard error of $\widehat{\textrm{ATC}}$ and subsequently constructed Wald-type confidence intervals.
4.5 Performance measures
To assess the performance of the estimators in our simulation study, we computed several key metrics: bias, empirical standard error (ESE), and 95% confidence interval (CI) coverage. Bias was calculated as the difference between the average of the point estimates across simulations and the true estimand value, providing a measure of systematic error. The ESE was computed as the standard deviation of the point estimates across simulations, reflecting the precision or variability of the different estimators. The 95% CI coverage was determined as the proportion of simulated datasets in which the constructed 95% confidence interval contained the true estimand value, evaluating the quality of interval estimation. In addition, we estimated Monte Carlo standard errors (MCSEs) using the formulas provided by Morris et al (2019) to quantify the uncertainty in the performance measures due to using a finite number of simulations.91
4.6 Results
The complete results for the simulation study are displayed in Tables 1-4. Note that the MCSEs are less than 0.001 for bias, ESE and 95% CI coverage, and considered negligible for all three performance metrics. Consequently, they have been omitted from Table 1 to Table 4.
As expected, the naïve estimator is biased in all four simulation scenarios. Under $n=1000$, the G-computation estimator is virtually unbiased in Scenario KS1 (bias = 0.003; see Table 1) and Scenario KS3 (bias = 0.004; see Table 3), where the outcome model is correctly specified; but exhibits bias where the outcome model is misspecified, in Scenario KS2 (bias = 0.067; see Table 2) and, particularly, in Scenario KS4 (bias = 0.516; see Table 4). Under $n=1000$, the IOW estimators based on modeling weights are virtually unbiased in Scenario KS1 and Scenario KS2, where the propensity score model for data source assignment is correctly specified; but are biased in Scenario KS3 and Scenario KS4, where the propensity score model is misspecified. Notably, in KS3 where the propensity score model is misspecified, the IOW estimator with normalized weights displays a bias in the opposite direction (and of a lesser magnitude) than the version with non-normalized weights.
While, under $n=1000$, the non-augmented MAIC (entropy balancing) estimator is unbiased in Scenarios KS1 and KS2, it exhibits bias in Scenarios KS3 and KS4, which confirms its lack of double robustness where the outcome-generating model is logistic. This motivates the need for explicit augmentation with an outcome model. All the augmented estimators appear to be unbiased with $n=1000$ where only one of either the propensity score model or the outcome model is correctly specified (Scenarios KS2 and KS3), but not both (Scenario KS4). This suggests that these estimators do have the doubly robust property and increased bias-robustness compared to MAIC, providing two opportunities for valid inference and additional protection against model misspecification.
The trends observed for the bias are similar under $n=200$ with the caveat that the (augmented and non-augmented) weighting estimators that appeared unbiased under $n=1000$, exhibit some small-sample bias in the corresponding scenarios. This is particularly notable in Scenario KS2 and, to a lesser extent, in Scenario KS1, and is probably due to small effective sample sizes after weighting.
Under correct specification of the outcome model (KS1 and KS3), G-computation is the most precise covariate-adjusted estimator, but the augmented estimators are almost as precise (e.g., compare the G-computation estimator which obtains ESE = 0.150 to the DR with MAIC weights estimator which obtains ESE = 0.169 for KS1 with $n=1000$; see Table 1), and generally produce precision gains versus their respective non-augmented weighting counterparts. When both the outcome model and the propensity score model are correctly specified (KS1), all augmented estimators have increased precision compared to the non-augmented weighting estimators based on modeling weights, but not necessarily against MAIC (any increase in precision for $n=1000$ is modest).
When only the outcome model is correctly specified (KS3), all augmented estimators display greater precision than the non-augmented weighting estimators, including MAIC. When only the propensity score model is correctly specified (KS2), outcome model misspecification does not induce a loss in precision for the augmented estimators compared to their non-augmented weighting counterparts, with the exception of IOW-based estimators with normalized modeling weights. All augmented estimators have similar precision to G-computation in KS2; albeit with $n=200$, the precision gains for G-computation are more notable, probably due to limited effective sample sizes after weighting.
As expected, when both the propensity score model and the outcome model are misspecified, all estimators are biased. Augmentation via an outcome model does not protect against the simultaneous misspecification of two models. There have been some concerns in the literature about doubly robust augmented estimators amplifying bias and variance when misspecified weights are combined with a misspecified outcome model.85 Such amplification does not take place in our simulation study; under dual model misspecification (KS4), the augmented estimators offer less bias and increased precision compared to their non-augmented weighting counterparts, and our proposed augmented MAIC estimator is the least biased of all estimators and the most precise of the augmented and non-augmented weighting estimators.
In Section 3.3, we hypothesized that entropy balancing weights, like those employed by MAIC, can lead to more stable and precise ATC estimation than inverse odds modeling weights. This appears to be confirmed for the non-augmented estimators in our simulation study; MAIC exhibits greater precision than the approaches using (normalized or non-normalized) IOW modeling weights in all simulation scenarios. Additionally, the precision gains have been inherited by the augmented approaches. For the methods highlighted in Section 3.4, estimators using MAIC weights display enhanced precision compared to those using IOW modeling weights in all simulation scenarios, while producing similar levels of bias, even lower bias under dual model misspecification.
Method Bias ESE 95% CI coverage Average 95% CI width
$n=200$
1. The naïve estimator 0.618 0.328 0.539 1.292
2. IOW with weights from modeling 0.024 0.528 0.939 1.978
3. IOW with normalized weights from modeling 0.049 0.456 0.944 1.763
4. MAIC 0.033 0.420 0.959 2.241
5. G-computation 0.016 0.350 0.955 1.430
6. DR with “modeling” IOW weights 0.029 0.421 0.954 1.667
7. DR with normalized “modeling” IOW weights 0.029 0.414 0.948 1.610
8. DR with MAIC weights 0.029 0.412 0.953 1.713
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.027 0.404 0.940 1.583
10. Augmented “weighted G-computation” with MAIC weights 0.026 0.406 0.943 1.740
$n=1000$
1. The naïve estimator 0.604 0.143 0.009 0.561
2. IOW with weights from modeling 0.009 0.205 0.950 0.806
3. IOW with normalized weights from modeling 0.010 0.196 0.941 0.750
4. MAIC 0.006 0.171 0.942 0.659
5. G-computation 0.003 0.150 0.949 0.592
6. DR with “modeling” IOW weights 0.005 0.174 0.946 0.675
7. DR with normalized “modeling” IOW weights 0.005 0.174 0.946 0.666
8. DR with MAIC weights 0.005 0.169 0.941 0.651
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.004 0.169 0.942 0.648
10. Augmented “weighted G-computation” with MAIC weights 0.004 0.169 0.940 0.649
Table 1: Results from Scenario KS1, where both the outcome model and the propensity score model are correctly specified.
Method Bias ESE 95% CI coverage Average 95% CI width
$n=200$
1. The naïve estimator 0.223 0.322 0.910 1.275
2. IOW with weights from modeling 0.022 0.665 0.929 2.386
3. IOW with normalized weights from modeling 0.052 0.515 0.938 1.954
4. MAIC 0.052 0.501 0.960 2.747
5. G-computation 0.081 0.436 0.951 1.731
6. DR with “modeling” IOW weights 0.043 0.542 0.947 2.100
7. DR with normalized “modeling” IOW weights 0.043 0.520 0.941 2.003
8. DR with MAIC weights 0.039 0.490 0.950 2.044
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.049 0.480 0.936 1.875
10. Augmented “weighted G-computation” with MAIC weights 0.028 0.480 0.946 2.198
$n=1000$
1. The naïve estimator 0.220 0.141 0.665 0.556
2. IOW with weights from modeling 0.012 0.277 0.946 1.077
3. IOW with normalized weights from modeling 0.007 0.221 0.938 0.837
4. MAIC 0.006 0.205 0.934 0.777
5. G-computation 0.067 0.188 0.936 0.734
6. DR with “modeling” IOW weights 0.005 0.226 0.941 0.865
7. DR with normalized “modeling” IOW weights 0.006 0.224 0.937 0.848
8. DR with MAIC weights 0.005 0.205 0.936 0.775
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.009 0.202 0.935 0.778
10. Augmented “weighted G-computation” with MAIC weights 0.005 0.203 0.935 0.769
Table 2: Results from Scenario KS2, where the outcome model is incorrectly specified and the propensity score model is correctly specified.
Method Bias ESE 95% CI coverage Average 95% CI width
$n=200$
1. The naïve estimator -0.033 0.301 0.952 1.208
2. IOW with weights from modeling 0.132 0.568 0.961 2.212
3. IOW with normalized weights from modeling -0.032 0.386 0.951 1.534
4. MAIC 0.121 0.346 0.955 1.518
5. G-computation 0.007 0.284 0.959 1.175
6. DR with “modeling” IOW weights 0.018 0.340 0.961 1.398
7. DR with normalized “modeling” IOW weights 0.017 0.328 0.957 1.335
8. DR with MAIC weights 0.015 0.310 0.956 1.290
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.012 0.302 0.954 1.242
10. Augmented “weighted G-computation” with MAIC weights 0.010 0.302 0.958 1.283
$n=1000$
1. The naïve estimator -0.040 0.134 0.937 0.528
2. IOW with weights from modeling 0.117 0.228 0.968 0.918
3. IOW with normalized weights from modeling -0.046 0.165 0.938 0.645
4. MAIC 0.104 0.146 0.889 0.573
5. G-computation 0.004 0.122 0.952 0.487
6. DR with “modeling” IOW weights 0.007 0.142 0.947 0.559
7. DR with normalized “modeling” IOW weights 0.007 0.140 0.946 0.549
8. DR with MAIC weights 0.006 0.132 0.946 0.517
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.005 0.127 0.949 0.505
10. Augmented “weighted G-computation” with MAIC weights 0.005 0.128 0.948 0.504
Table 3: Results from Scenario KS3, where the outcome model is correctly specified and the propensity score model is incorrectly specified.
In Section 3.5, we described an augmented “weighted G-computation” estimator that is doubly robust where the target estimand is the ATE and the outcome model is a GLM with a canonical link function. The results of the simulation study suggest that this estimator is also doubly robust for the ATC, noting that the employed logistic outcome model has a canonical link function (the logit link for a binomial outcome distribution). Interestingly, this augmented “weighted G-computation” estimator offers very similar performance, both in terms of bias and precision, than our proposed doubly robust augmented estimator with MAIC weights, regardless of whether (normalized) modeling IOW weights or MAIC weights are used. These three estimators are generally the least biased and most precise of all the augmented and non-augmented weighting estimators. Their similar performance is consistent with theoretical investigations by Sloczynski et al (2023),93 which suggest that the different versions of augmented weighting estimators are numerically equivalent when balancing weights are used and the target estimand is the ATT or the ATC.aaaIt is worth noting that the augmented weighting estimators investigated by Sloczynski et al (2023) use a different type of balancing weights, those from Imai and Ratkovic (2014),94 and a linear outcome model.93
Method Bias ESE 95% CI coverage Average 95% CI width
$n=200$
1. The naïve estimator 0.519 0.338 0.684 1.329
2. IOW with weights from modeling 0.800 0.783 0.908 2.763
3. IOW with normalized weights from modeling 0.573 0.475 0.757 1.832
4. MAIC 0.608 0.469 0.787 2.013
5. G-computation 0.532 0.383 0.745 1.544
6. DR with “modeling” IOW weights 0.576 0.459 0.767 1.831
7. DR with normalized “modeling” IOW weights 0.571 0.443 0.750 1.753
8. DR with MAIC weights 0.513 0.414 0.774 1.664
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.541 0.420 0.761 1.678
10. Augmented “weighted G-computation” with MAIC weights 0.540 0.429 0.781 1.780
$n=1000$
1. The naïve estimator 0.497 0.146 0.071 0.574
2. IOW with weights from modeling 0.788 0.359 0.334 1.425
3. IOW with normalized weights from modeling 0.520 0.199 0.249 0.765
4. MAIC 0.552 0.192 0.165 0.738
5. G-computation 0.516 0.162 0.104 0.635
6. DR with “modeling” IOW weights 0.542 0.188 0.178 0.729
7. DR with normalized “modeling” IOW weights 0.541 0.185 0.170 0.716
8. DR with MAIC weights 0.487 0.173 0.188 0.665
9. Augmented “weighted G-computation” with normalized “modeling” IOW weights 0.530 0.178 0.148 0.710
10. Augmented “weighted G-computation” with MAIC weights 0.539 0.183 0.153 0.708
Table 4: Results from Scenario KS4, where both the outcome model and the propensity score model are incorrectly specified.
Assuming unbiasedness, interval estimation is appropriate if the coverage is approximately equal to 0.95; poor coverage can arise due to bias or to inadequate variance/interval estimation. Coverage is generally close to 0.95 for all covariate adjustment methods in the simulation scenarios under which they are unbiased, which suggests that our proposed non-parametric bootstrap approach for variance estimation is adequate. In the cases in which the covariate adjustment methods are unbiased, there is slight undercoverage for $n=1000$, with coverage rates between 0.934 and 0.950 across scenarios. Interestingly, coverage rates seem to increase for these methods under $n=200$ despite the small-sample bias, lying between 0.938 and 0.961 (except IOW with modeling weights in KS2 with 0.929), probably due to overly wide interval estimates due to small sample sizes. Note that, due to computational limitations, the non-parametric bootstrap approach in the simulation study was conducted with only $B=100$ resamples. This may have impacted the observed coverage rates and we suspect that coverage might be more appropriate when using a larger number of resamples.
While some covariate adjustment methods display bias-induced undercoverage in the scenarios under which they are biased (e.g., MAIC in KS3 under $n=1000$ or all estimators in KS4), they may also display adequate coverage (e.g., IOW with normalized modeling weights under $n=200$) because of excessively large standard errors, probably due to low effective sample sizes after weighting. As observed for KS4, bias-induced undercoverage tends to worsen with higher sample sizes, as interval estimates around the wrong target value become narrower. The naïve estimator displays discernible undercoverage in KS1, KS2 and KS4 (particularly under $n=1000$), not only due to bias but also due to overprecise standard errors that do not account for covariate differences.
5 Applied example
We now demonstrate the application of some of the methods outlined in Section 3 to synthetic lung cancer clinical trial data. The data were obtained from the “MAIC” R package, implemented by researchers from the pharmaceutical industry.95 Our objective is to compare the objective response, a binary outcome $Y$, under two treatments: the active “intervention” ($T=1$) and the external “control” ($T=0$). The data consist of IPD from a SAT ($S=1$) with $n_{1}=500$ subjects under the “intervention”, and AD from an external historical SAT ($S=0$) of $n_{0}=300$, which makes up the “control”. The unavailability of IPD for the external control allows us to illustrate the methodological extensions described in Section 3.7. R code to reproduce the applied example is provided in the Supplementary Material.
The target estimand is the ATC on the marginal log-odds ratio scale. Four baseline covariates, one continuous – age – and three binary – sex, the Eastern Cooperative Oncology Group (ECOG) performance status and smoking status — have been identified as prognostic factors under the intervention, and are imbalanced between the intervention SAT and external control samples. There are no missing values for baseline characteristics and outcomes. Subjects in the intervention SAT are, on average, somewhat older, less likely to be male, more likely to be physically restricted (as indicated by ECOG performance status), and more likely to be smokers, relative to subjects in the external control (Table 5). In addition, the age of subjects in the intervention SAT has substantially greater variance than that of subjects in the external control.
Covariate Intervention SAT External control Weighted intervention SAT
[…]
Age in years 59.85; 9.01 50.06; 3.24 50.06; 3.24
(mean; standard deviation)
Sex 0.38 0.49 0.49
(proportion male)
ECOG 0.41 0.35 0.35
(proportion ECOG performance status of 1)
Smoking status 0.32 0.19 0.19
(proportion of smokers)
Table 5: Summary statistics of the four baseline covariates identified as imbalanced prognostic factors. Note that the standard deviation of age in the weighted intervention SAT is $\sum_{i=1}^{n_{1}}\left(\omega_{i}(X_{1,i}-\sum_{i=1}^{n_{1}}\omega_{i}X_{1,i})^{2}\right)$, where $X_{1,i}$ and $\omega_{i}$ are the age and the MAIC weight, respectively, for subject $i=1,\dots,n_{1}$ in the intervention SAT.
We consider the naïve estimator first, which does not perform covariate adjustment (Equation 35). In the intervention SAT, 390 of $n_{1}=500$ subjects attained objective response, which equates to a 78% response rate, $\hat{\mu}_{1}^{1}=0.78$. This implies a log-odds of response of $g(\hat{\mu}_{1}^{1})=1.266$, where $g(\cdot)=\operatorname{logit}(\cdot)$. In the external control, 120 of $n_{0}=300$ subjects attained objective response, which equates to a 40% response rate, $\hat{\mu}_{0}^{0}=0.4$. This implies a log-odds of response of $g(\hat{\mu}_{0}^{0})=-0.405$. A naïve estimate is obtained by simple subtraction:
$\widehat{\textrm{ATC}}_{naive}=1.266-(-0.405)=1.671.$
Using the Delta method, we obtain:
$\textrm{SE}\left(g(\hat{\mu}_{1}^{1})\right)$ $=\sqrt{{1}/{500(0.78\times(1-0.78))}}$
$=0.108,$
$\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)$ $=\sqrt{{1}/{300(0.40\times(1-0.40))}}$
$=0.118.$
Then, following Equation 33, we have:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{naive}\right)$ $=\sqrt{(\textrm{SE}\left(g\left(\hat{\mu}_{1}^{1})\right)\right)^{2}+(\textrm{SE}\left(g\left(\hat{\mu}_{0}^{0})\right)\right)^{2}}$
$=\sqrt{0.108^{2}+0.118^{2}}$
$=0.160.$
Based on the computed standard errors, Wald-type 95% confidence intervals are given by:
$(\widehat{\textrm{ATC}}_{naive}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{naive}),\widehat{\textrm{ATC}}_{naive}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{naive}))=(1.358,1.984)$.
We now consider the MAIC estimator. The positivity assumption is assessed using a method proposed by Glimm and Yau (2022), which verifies whether covariate AD from the external control lie within the convex hull of the SAT covariate space, and whether the MAIC numerical optimization algorithm can converge.60 This method is implemented using the “maicLP” R function in the “maicChecks” R package,96 which confirms that a feasible weighting solution to the MAIC convex optimization problem exists, i.e., that there is a set of positive weights that can enforce covariate balance between the intervention SAT and the external control, and that the MAIC numerical optimization algorithm can converge.
MAIC is performed using the procedure described in Section 3.3. We choose to weight the intervention SAT such that the means of all four baseline covariates and the variance of age are exactly balanced with respect to the external control. Following the notation in Section 3.3, we have:
$\mathbf{c(X)}=[\mathbf{Age},\mathbf{Sex},\mathbf{ECOG},\mathbf{Smoking},\mathbf{Age}^{2}]^{\top}=\begin{pmatrix}45&71&\cdots&58\\ 1&1&\cdots&0\\ 0&0&\cdots&1\\ 0&0&\cdots&1\\ 2025&5041&\cdots&3364\end{pmatrix},$
where $\mathbf{c(X)}$ is a 5-by-500 matrix, with the rows representing the age, sex, ECOG performance status, smoking status, and age-squared for subjects in the intervention SAT. Using the BFGS convex optimization algorithm to minimize the objective function in Equation 13, we obtain $\mathbf{\hat{\gamma}}=(3.542,0.589,-0.698,-0.048,-0.036)$, and weights are calculated subject to the constraint that they sum to one. Figure 1 shows a histogram illustrating the empirical distribution of the resulting MAIC weights. The effective sample size (ESS) – that is, the number of independent non-weighted observations that would be required to give an estimate with approximately the same precision as the weighted sample estimate – of the intervention SAT after weighting is:
$\textrm{ESS}=\frac{\left(\sum_{i=1}^{n_{1}}\omega_{i}\right)^{2}}{\sum_{i=1}^{n_{1}}\left(\omega_{i}^{2}\right)}=\frac{1}{0.0063665}=157.07.$
Figure 1: Histogram of the EB (MAIC) weights; EB denotes entropy balancing.
Using the MAIC weights, the ATC is estimated as:
$\widehat{\textrm{ATC}}_{MAIC}=$ $g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-\hat{\mu}_{0}^{0}$
$=0.926-(-0.405)$
$=1.331.$
We previously calculated $\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)=0.118$. Using the non-parametric bootstrap with $B=10000$ resamples, $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.177$. As per Equation 33:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{MAIC}\right)$ $=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}}$
$=\sqrt{0.177^{2}+0.118^{2}}$
$=0.212,$
and Wald-type 95% confidence intervals are given by
$(\widehat{\textrm{ATC}}_{MAIC}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{MAIC}),\widehat{\textrm{ATC}}_{MAIC}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{MAIC}))=(0.915,1.748)$.
To perform G-computation with unavailable IPD for the external control, we first simulate $M=10000$ individual-level covariate profiles from an assumed joint covariate distribution for the external control, as per the approach outlined by Remiro-Azócar et al (2022)46 or the “infinite population” STC method described by Zhang et al (2024).97 We proceed by assuming that the pairwise correlation structure of the four covariates in the external control is equal to that observed in the intervention SAT:
$\begin{pmatrix}1.00&0.03&0.00&0.00\\ 0.03&1.00&-0.14&-0.02\\ 0.00&-0.14&1.00&-0.01\\ 0.00&-0.02&-0.01&1.00\\ \end{pmatrix},$
with the rows/columns in the order: age, sex, ECOG performance status and smoking status. For age, we assumed a Normal(50.06, 3.24) marginal distribution; for sex, ECOG performance status and smoking status, we assumed Bernoulli(0.49), Bernoulli(0.35) and Bernoulli(0.19) marginal distributions, respectively, based on the summary statistics of the external control in Table 5. Individual-level covariates were ultimately simulated from a Gaussian copula characterized by the aforementioned marginal distributions and pairwise correlation structure, using the “add_integration” function from the multinma R package.98
Subsequently, a logistic-link binomial GLM for the outcome expectation under the intervention, conditional on baseline covariates, was postulated. This relates objective response $Y_{i}^{1}$ under the intervention $T=1$ to baseline covariates $\mathbf{X}_{i}=\left(Age_{i},Sex_{i},ECOG_{i},Smoking_{i}\right)^{\top}$ as:
$\textrm{logit}\left(\textrm{E}(Y_{i}^{1}\mid\mathbf{X}_{i};\mathbf{\beta})\right)=\beta_{0}+\beta_{1}Age_{i}+\beta_{2}Sex_{i}+\beta_{3}ECOG_{i}+\beta_{4}Smoking_{i}+\beta_{5}Age_{i}^{2},$
for $i$ in $n_{1}$, where $\mathbf{\beta}=\left(\beta_{0},\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5}\right)^{\top}$ are regression coefficients. The model was fitted to the intervention SAT using maximum-likelihood estimation with Fisher scoring, and we obtained regression coefficient estimates of $\hat{\beta}_{0}=5.72$, $\hat{\beta}_{1}=-0.20$, $\hat{\beta}_{2}=0.12$, $\hat{\beta}_{3}=0.13$, $\hat{\beta}_{4}=0.01$, and $\hat{\beta}_{5}=0.00$.
Then, following the G-computation procedure outlined in Section 3.4, the ATC is estimated as:
$\widehat{\textrm{ATC}}_{Gcomp}$ $=g\underbrace{\left(\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y}_{i}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right)$
$=g{\left(\frac{7149.896}{10000}\right)}-(-0.405)$
$=1.325,$
where $\hat{Y}_{i}^{1}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$ for $i=n_{1}$,…,$n_{1}+M$, are predictions of the fitted outcome model – indexed $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})$ with regression coefficient point estimates $\hat{\mathbf{\beta}}$ – for each of the $M$ simulated individual-level covariate profiles.
Using the non-parametric bootstrap with $B=10000$ resamples, $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.164$. As per Equation 33:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{Gcomp}\right)$ $=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}}$
$=\sqrt{0.164^{2}+0.118^{2}}$
$=0.202,$
and Wald-type 95% confidence intervals are given by
$(\widehat{\textrm{ATC}}_{Gcomp}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{Gcomp}),\widehat{\textrm{ATC}}_{Gcomp}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{Gcomp}))=(0.929,1.722)$.
Finally, our doubly robust augmented MAIC estimator proposed in Section 3.4 produces the estimate:
[…]
$=g\underbrace{\left(0.001+\frac{7149.896}{10000}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right)$
$=0.926-(-0.405)$
$=1.332,$
where $\hat{Y}_{i}^{1}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$ are predictions of the fitted outcome model for each of the subjects in the intervention SAT ($i=1,...,n_{1}$) and each of the $M$ simulated individual-level covariate profiles ($i=n+1,\dots,n_{1}+M$).
Using the non-parametric bootstrap with $B=10000$ resamples, $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.179$. As per Equation 33:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{DR}\right)$ $=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}}$
$=\sqrt{0.179^{2}+0.118^{2}}$
$=0.214,$
and Wald-type 95% confidence intervals are given by
$(\widehat{\textrm{ATC}}_{DR}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{DR}),\widehat{\textrm{ATC}}_{DR}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{DR}))=(0.912,1.751)$.
Figure 2 shows the point estimates obtained using the four different estimators alongside their 95% confidence intervals. When comparing the covariate-adjusted approaches to the naïve approach, we observe that covariate adjustment shifts the point estimate towards the null considerably.
Figure 2: Point estimates with 95% confidence intervals of the ATC (marginal log-odds ratio of objective response) for the different estimators in the applied example. DR denotes doubly robust and EB denotes entropy balancing.
When comparing MAIC, G-computation and our proposed doubly robust (DR) augmented MAIC estimator, results across the three methods seem consistent. Despite the shift towards the null, results suggest that the intervention improves objective response versus the control, statistically significant at the 5% level. In this case, the DR point estimate is not meaningfully different than the MAIC or G-computation point estimates, and the DR approach offers slightly increased standard errors and wider confidence intervals than G-computation. Nevertheless, this loss of precision seems a relatively minor price to pay, compensated for by greater reassurance in our results due to increased protection against misspecification of the outcome model.
6 Discussion
The objective of this article was to clarify existing approaches for doubly robust estimation in the context of externally controlled SATs, and to propose a novel estimator that augments approaches based on MAIC or entropy balancing via an outcome model. We described and illustrated an extension of this estimator specifically tailored to unanchored ITCs, for the setting with unavailable external control IPD, which is commonly encountered in practice. In a simulation study and applied example, we evaluated the performance and demonstrated the use of different doubly robust augmented estimators, highlighting their merits with respect to the more popular non-augmented singly robust estimators.
Our findings reinforce the understanding that “balancing” approaches to weighting, such as MAIC or entropy balancing, can enhance performance relative to standard “modeling” approaches, but lack the doubly robust property for non-linear outcome models. Conversely, the augmented MAIC or entropy balancing-based estimator demonstrates double robustness with a logistic outcome model, and exhibits higher precision than non-augmented weighting estimators when the outcome model is correctly specified. Moreover, it generally achieves near-identical precision to G-computation, which offers the lowest variance under correct specification of the outcome model, but may exhibit notable bias where the outcome model is misspecified.
A potential concern about augmented estimators has been possible bias and variance amplification where both the propensity score and the outcome model are incorrectly specified. Nevertheless, such amplification under dual model misspecification was not observed in our simulation study, and one can argue that risks are mitigated by employing “balancing” instead of “modeling” weights (for reasons outlined in the introduction to Section 3.3), but further theoretical work and simulation studies are required to fully support this claim. The results of our simulation study motivate the routine application of doubly robust augmented estimators, particularly those based on MAIC or entropy balancing weights, in practical settings. It is unfortunate to see that virtually all applications of MAIC do not consider augmentation via an outcome model, and that most practical uses of augmented estimators apply “modeling” instead of “balancing” weights.99
While the “weighted G-computation” estimator described in Section 3.5 exhibited double robustness for the ATC in our simulation study and almost identical performance to our proposed augmented MAIC estimator, these results are only expected to hold where the outcome model is a GLM with canonical link function, and not where the link function is non-canonical or for outcome models in the time-to-event setting.51 Future simulation studies should consider non-binary outcomes, including survival outcomes with censoring, and other summary effect measures beyond the log-odds ratio.
Our simulation study and applied example considered scenarios with relatively low sample sizes in the SAT and the external control, corresponding to typical settings in rare disease and late-stage hematological or solid tumor oncology, where the number of subjects enrolled in SATs can be one- or two-hundred, but may also consist of several hundreds. The number of external controls can be equally small; as such, our findings are potentially applicable where the target estimand is the ATT instead the ATC, which requires weighting the external control as opposed to the SAT. Somewhat worryingly, all augmented and non-augmented weighting estimators displayed some small-sample bias in our simulation study under a total sample size of $n=200$, even if modeling assumptions were correct.
The level of (deterministic) overlap between the SAT and external control covariate distributions in our simulation study was relatively strong. The performance of augmented and non-augmented weighting estimators with respect to G-computation will likely worsen as overlap decreases, particularly in conjunction with small sample sizes. Nevertheless, we conjecture that the performance of balancing-based approaches will suffer less than that of their corresponding modeling-based counterparts, due to generating more stable and less extreme weights. This is unless a complete lack of overlap results in the absence of a solution to the convex optimization balancing problem, in which case extrapolating via G-computation would be the only option.
Finally, it is important to emphasize that all the covariate-adjusted estimators we have considered, including the doubly robust augmented approaches, require the important assumption of no unmeasured prognostic factors, and can incur bias if these are omitted or missing. In practice, important prognostic factors may be unknown or unavailable in at least one of the SAT or external control data sources. An important area of future research is the development of sensitivity analysis or quantitative bias analysis methods to help explore the sensitivity of results to unmeasured prognostic factors, in the specific context of externally controlled SATs and unanchored ITCs.100
Acknowledgments
Not available.
Financial disclosure
No funding to report.
Conflict of interest
Harlan Campbell is employed by Precision AQ, a life sciences consultancy company, and Antonio Remiro-Azócar is employed by Novo Nordisk, a pharmaceutical company. No conflicts of interest are declared as this research is purely methodological.
Data Availability Statement
The files required to generate the data, run the simulations, and reproduce the results of the simulation study are available at https://github.com/harlanhappydog/DRAWE-. R code to reproduce the applied example is provided in the Supplementary Material.
Supplementary Material
Plots of covariate overlap for the simulation study
Figure 3: Density plots showing the overlap of covariates $X_{1}$,$X_{2}$,$X_{3}$, and $X_{4}$ for Scenarios KS1 and KS2 in the simulation study.
Figure 4: Density plots showing the overlap of covariates $X_{1}$,$X_{2}$,$X_{3}$, and $X_{4}$ for Scenarios KS3 and KS4 in the simulation study.
R code for the applied example
library(dplyr)
library(boot)
library(survival)
library(MAIC)
library(ggplot2)
library(survminer)
library(flextable)
library(officer)
# set seed to give reproducible example
set.seed(1894)
boot_n<-10000
# g-function is the log-odds:
g_function <- function(p){log(p/(1-p))}
#### Intervention data
# Read in ADaM data and rename variables of interest
adsl <- read.csv(system.file("extdata", "adsl.csv", package = "MAIC", mustWork = TRUE))
adrs <- read.csv(system.file("extdata", "adrs.csv", package = "MAIC", mustWork = TRUE))
adtte <- read.csv(system.file("extdata", "adtte.csv", package = "MAIC", mustWork = TRUE))
adsl <- adsl %># Data containing the matching variables
  mutate(SEX=ifelse(SEX=="Male", 1, 0)) # Coded 1 for males and 0 for females
adrs <- adrs %># Response data
  filter(PARAM=="Response") %>%
  transmute(USUBJID, ARM, response=AVAL)
adtte <- adtte %># Time to event data (overall survival)
  filter(PARAMCD=="OS") %>%
  mutate(Event=1-CNSR) %>#Set up coding as Event = 1, Censor = 0
  transmute(USUBJID, ARM, Time=AVAL, Event)
# Combine all intervention data
intervention_input <- adsl %>%
  full_join(adrs, by=c("USUBJID", "ARM")) %>%
  full_join(adtte, by=c("USUBJID", "ARM"))
head(intervention_input)
# List out matching covariates
match_cov <- c("AGE",
               "SEX",
               "SMOKE",
               "ECOG0")
# Baseline aggregate data for the comparator population
target_pop <- read.csv(system.file("extdata", "aggregate_data.csv",
                                   package = "MAIC", mustWork = TRUE))
# Renames target population cols to be consistent with match_cov
match_cov
names(target_pop)
target_pop_standard <- target_pop %>%
  #EDIT
  dplyr::rename(N=N,
         Treatment=ARM,
         AGE=age.mean,
         SEX=prop.male,
         SMOKE=prop.smoke,
         ECOG0=prop.ecog0
  transmute(N, Treatment, AGE, SEX, SMOKE, ECOG0)
target_pop_standard
# Simulate response data based on the known proportion of responders
comparator_n <- target_pop\$N # total number of patients in the comparator data
comparator_prop_events <- 0.4 # proportion of responders
# Calculate number with event
# Use round() to ensure we end up with a whole number of people
# number without an event = Total N - number with event to ensure we keep the same number of patients
n_with_event <- round(comparator_n*comparator_prop_events, digits = 0)
comparator_binary <- data.frame("response"= c(rep(1, n_with_event), rep(0, comparator_n - n_with_event)))
n0 <- dim(comparator_binary)[1]
n1 <- length(intervention_input\$response)
Y_all <- unlist(c(intervention_input\$response,comparator_binary))
S_all <- c(rep(1, n1), rep(0,n0))
X_all <- rbind(cbind(intervention_input%>%select(AGE, SEX, SMOKE, ECOG0)),
               cbind(AGE=rep(NA, n0),
                     SEX=rep(NA, n0),
                     SMOKE=rep(NA, n0),
                     ECOG0=rep(NA, n0)))
X_all\$AGE_SQ<-(X_all\$AGE)^2
  n1 <- sum(S_all==1)
  n <- n1 + n0
  n1/n0
  # naive estimate
  mu1_naive <- mean(Y_all[S_all==1])
  mu0_naive <- mean(Y_all[S_all==0])
  ATC_naive =  g_function(mu1_naive) - g_function(mu0_naive)
  ATC_naive
  SE_g_mu1 <- sqrt(1/(length((Y_all[S_all==1]))*mu1_naive*(1-mu1_naive)))
  SE_g_mu0 <- sqrt(1/(length((Y_all[S_all==0]))*mu0_naive*(1-mu0_naive)))
  (sqrt((mu0_naive*(1-mu0_naive)/length((Y_all[S_all==0])))))
  naive_function <- function(data, indices) {
    temp <- data[indices,]
    return( g_function(mean(temp[temp[,"S_all"]==1, "Y_all"])))}
  boot_samples <- boot(data=data.frame(Y_all=Y_all, S_all=S_all),
                       statistic=naive_function, R=boot_n,
                       strata=S_all,
                       parallel = "multicore")
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  SE_ATC
  ATC_naive_CI_boot<- c(ATC_naive-abs(qnorm(0.025))*SE_ATC, ATC_naive+abs(qnorm(0.025))*SE_ATC)
  round(ATC_naive_CI_boot,3)
  SE_ATC <- sqrt(SE_g_mu1^2  + SE_g_mu0^2)
  ATC_naive_CI<- c(ATC_naive-abs(qnorm(0.025))*SE_ATC, ATC_naive+abs(qnorm(0.025))*SE_ATC)
  round(ATC_naive_CI,3)
  round((c(ATC_naive, ATC_naive_CI)),3)
  round(exp(c(ATC_naive, ATC_naive_CI)),3)
  # 5.318 3.880 7.289
  # compare with published result:
  # 5.318 (3.888 to 7.275)  https://roche.github.io/MAIC/articles/MAIC.html
  # 3.3 Entropy balancing (matching-adjusted indirect comparison)
  # ...equivalent to minimizing the objective function:
  objfn <- function(a1, X){ sum(exp(X %*a1)) }
  gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*a1), "*")) }
  cov_names <- paste0("X.", colnames(X_all))
  AC.IPD <- data.frame(y=Y_all[S_all%in%c(1)], X= X_all[S_all==1,])
  BC.ALD <- data.frame(cbind(target_pop\$age.mean,
                         target_pop\$prop.male,
                         target_pop\$prop.smoke,
                         target_pop\$prop.ecog0,
                         target_pop\$age.mean^2 + target_pop\$age.sd^2))
  colnames(BC.ALD)<- paste0("mean.",cov_names)
  library(maicChecks)
  maicLP(AC.IPD[,-1], BC.ALD) # Checks if AD is within the convex hull of IPD using lp-solve
  X.EM.0 <- sweep(cbind(as.matrix((AC.IPD[, cov_names]))), 2,
                  as.matrix((BC.ALD[,c(paste("mean.",cov_names, sep=""))])), ’-’)
  gamma <- optim(par = rep(0,dim(X.EM.0)[2]),
                 fn = objfn, gr = gradfn, X = X.EM.0, method = "BFGS")\$par
  wt_EB <- exp(X.EM.0 %*gamma)/sum(exp(X.EM.0 %*gamma))
  mu1_EB <- sum(wt_EB*Y_all[S_all==1])
  mu0_EB <- mu0_naive
  ATC_EB <- g_function(mu1_EB) - g_function(mu0_EB)
  ATC_EB
  hist(wt_EB, xlab="Entropy balancing weights", main="")
  ATC_EB_function <- function(data, indices){
    temp <- data[indices,]
    AC.IPD <- data.frame(y=temp[,"Y_all"][temp[,"S_all"]==1],
                         X= temp[temp[,"S_all"]==1,grep("X_all",colnames(temp))])
    BC.ALD <- data.frame(cbind(target_pop\$age.mean,
                               target_pop\$prop.male,
                               target_pop\$prop.smoke,
                               target_pop\$prop.ecog0,
                               target_pop\$age.mean^2 + target_pop\$age.sd^2))
    objfn <- function(a1, X){ sum(exp(X %*a1)) }
    gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*a1), "*")) }
    cov_names <- paste0("X.X_all.", colnames(X_all))
    colnames(BC.ALD)<- paste0("mean.",cov_names)
    X.EM.0 <- sweep(cbind(as.matrix((AC.IPD[, cov_names]))), 2,
                    as.matrix((BC.ALD[,c(paste("mean.",cov_names, sep=""))])), ’-’)
    gamma <- optim(par = rep(0,dim(X.EM.0)[2]),
                   fn = objfn, gr = gradfn, X = X.EM.0, method = "BFGS")\$par
    wt_EB <- exp(X.EM.0 %*gamma)/sum(exp(X.EM.0 %*gamma))
    mu1_EB <- sum(wt_EB*temp[,"Y_all"][temp[,"S_all"]==1])
    mu0_naive <- mean(temp[temp[,"S_all"]==0,"Y_all"])
    return(g_function(mu1_EB))
  set.seed(123)
  boot_samples <- boot(data=data.frame(Y_all=Y_all, S_all=S_all, X_all=X_all),
                       statistic=ATC_EB_function, R=boot_n,
                       strata=S_all, parallel = "multicore")
  sd(boot_samples\$t,na.rm=TRUE)
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  SE_ATC
  ATC_EB_CI <- c(ATC_EB-abs(qnorm(0.025))*SE_ATC,ATC_EB+abs(qnorm(0.025))*SE_ATC)
  round((c(ATC_EB, ATC_EB_CI)),3)
  round(exp(c(ATC_EB, ATC_EB_CI)),3)
  # 3.787 2.497 5.742
  # compare with published result:
  # 3.787 (2.558 to 5.605)  https://roche.github.io/MAIC/articles/MAIC.html
  # Simulate M individual values from target population
  M = 10000
  set.seed(123)
  out2<- add_integration(data.frame(Y_all=NA),
                         AGE = distr(qnorm, mean=target_pop\$age.mean, sd=target_pop\$age.sd),
                         SEX = distr(qbern, prob=target_pop\$prop.male),
                         SMOKE = distr(qbern, prob=target_pop\$prop.smoke),
                         ECOG0 = distr(qbern, prob=target_pop\$prop.ecog0),
                         cor =  cor(X_all[S_all==1,c("AGE", "SEX", "SMOKE", "ECOG0")]),
                         n_int = M)
  x_star <-  cbind(unlist(out2\$.int_AGE),
                   unlist(out2\$.int_SEX),
                   unlist(out2\$.int_SMOKE),
                   unlist(out2\$.int_ECOG0))
  dim((X_all[S_all==1,]))
  dim((X_all[S_all==0,]))
  dim(x_star)
  # add squared age values:
  x_star <- cbind(x_star,(x_star[,1])^2)
  dim(x_star)
  #add names
  colnames(x_star)<-colnames(X_all[S_all==1,])
  n_with_event <- round(M*comparator_prop_events, digits = 0)
  Y_all <-  c(Y_all[S_all==1], c(rep(1, n_with_event), rep(0, M - n_with_event)))
  X_all <- rbind(X_all[S_all==1,], x_star)
  S_all <-  c(S_all[S_all==1], rep(0,M))
  dim((X_all[S_all==1,]))
  dim((X_all[S_all==0,]))
  n1 <- sum(S_all==1)
  n0 <- sum(S_all==0)
  n <- n1 + n0
  # The G-computation estimator
  # The G-computation estimator for the ATC contrasts the average of potential counterfactual outcomes
  # under the active intervention with the average of observed outcomes for the external control.
  outcome_model <- glm(y ~ .,
                       data = data.frame(y=c(Y_all[S_all==1]),
                                         x=(X_all[S_all==1,])),
                       family = binomial(link="logit"))Ψ
  Y1_hat <- (predict(outcome_model, newdata = data.frame(x=X_all[S_all==0,]),
                     type = "response"))
  mu1_GCOMP <- (1/n0)*sum(Y1_hat)
  mu0_GCOMP1 <- mu0_naive
  ATC_GCOMP <-  g_function(mu1_GCOMP) - g_function(mu0_GCOMP1)
  ATC_GCOMP
  ATC_GCOMP_function <- function(data, indices){
    temp <- data[indices,]
    outcome_model <- glm(y ~ .,
                         data = data.frame(y=c(temp[,"Y_all"][temp[,"S_all"]==1]),
                                           x=temp[temp[,"S_all"]==1,grep("X_all",colnames(temp))]),
                         family = "binomial")Ψ
    Y1_hat <- (predict(outcome_model, newdata = data.frame(x=temp[temp[,"S_all"]==0,grep("X_all",colnames(temp))]),
                       type = "response"))
    mu1_GCOMP <- (1/sum(temp[,"S_all"]==0))*sum(Y1_hat)
    return(g_function(mu1_GCOMP))}
  set.seed(123)
  boot_samples <- boot(data=data.frame(Y_all=Y_all, S_all=S_all, X_all=X_all),
                       statistic=ATC_GCOMP_function, R=boot_n,
                       strata=S_all, parallel = "multicore")
  sd(boot_samples\$t,na.rm=TRUE)
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  ATC_GCOMP_CI <- c(ATC_GCOMP-abs(qnorm(0.025))*SE_ATC,ATC_GCOMP+abs(qnorm(0.025))*SE_ATC)
  round((c(ATC_GCOMP, ATC_GCOMP_CI)),3)
  round(exp(c(ATC_GCOMP, ATC_GCOMP_CI)),3)
  # DR3
  # Our novel contribution is augmenting the entropy balancing-based MAIC approach with an outcome model
  data_for_outcome_model <- data.frame(y=c(Y_all[S_all==1]),
                                       X_all[S_all==1,])
  colnames(data_for_outcome_model)<-c("y", colnames(X_all))
  outcome_model <- glm(y ~ .,
                       data = data_for_outcome_model,
                       family = "binomial")Ψ
  Y1_hat_all <- (predict(outcome_model, newdata = data.frame(X_all),
                         type = "response"))
  mu1_DR3 <- (1/sum(wt_EB))*sum(wt_EB*(Y_all[S_all==1]  - Y1_hat_all[S_all==1])) +
    (1/n0)*sum(Y1_hat_all[S_all==0])
  mu0_DR3 <- mu0_naive
  ATC_DR3 <-  g_function(mu1_DR3) - g_function(mu0_DR3)
  ATC_DR3
  ATC_DR3_function <- function(data, indices){
    temp <- data[indices,]
    outcome_model <- glm(y ~ .,
                         data = data.frame(y=c(temp[,"Y_all"][temp[,"S_all"]==1]),
                                           x=temp[temp[,"S_all"]==1,grep("X_all",colnames(temp))]),
                         family = "binomial")Ψ
    Y1_hat_all <- (predict(outcome_model, newdata = data.frame(x=temp[,grep("X_all",colnames(temp))]),
                           type = "response"))
    AC.IPD <- data.frame(y=temp[,"Y_all"][temp[,"S_all"]==1], X= temp[temp[,"S_all"]==1,grep("X_all",colnames(temp))])
    BC.ALD <- data.frame(matrix(apply(temp[temp[,"S_all"]==0,grep("X_all",colnames(temp))],2,mean),1,))
    objfn <- function(a1, X){ sum(exp(X %*a1)) }
    gradfn <- function(a1, X){ colSums(sweep(X, 1, exp(X %*a1), "*")) }
    cov_names <- paste0("X.X_all.", colnames(X_all))
    colnames(BC.ALD)<- paste0("mean.",cov_names)
    X.EM.0 <- sweep(cbind(as.matrix((AC.IPD[, cov_names]))), 2,
                    as.matrix((BC.ALD[,c(paste("mean.",cov_names, sep=""))])), ’-’)
    gamma <- optim(par = rep(0,dim(X.EM.0)[2]),
                   fn = objfn, gr = gradfn, X = X.EM.0, method = "BFGS")\$par
    wt_EB <- exp(X.EM.0 %*gamma)/sum(exp(X.EM.0 %*gamma))
    mu1_DR3 <- (1/sum(wt_EB))*sum(wt_EB*(temp[,"Y_all"][temp[,"S_all"]==1]  -
                                           Y1_hat_all[temp[,"S_all"]==1])) +
      (1/sum(temp[,"S_all"]==0))*sum(Y1_hat_all[temp[,"S_all"]==0])
    return( g_function(mu1_DR3))}
  set.seed(123)
  boot_samples <- boot(data=data.frame(Y_all=Y_all, S_all=S_all, X_all=X_all),
                       statistic=ATC_DR3_function, R=boot_n,
                       strata=S_all, parallel = "multicore")
  sd(boot_samples\$t,na.rm=TRUE)
  SE_ATC <- sqrt(sd(boot_samples\$t,na.rm=TRUE)^2  + SE_g_mu0^2)
  SE_ATC
  ATC_DR3_CI <- c(ATC_DR3-abs(qnorm(0.025))*SE_ATC,ATC_DR3+abs(qnorm(0.025))*SE_ATC)
  ATC_DR3_CI
  round((c(ATC_DR3, ATC_DR3_CI)),3)
  round(exp(c(ATC_DR3, ATC_DR3_CI)),3)